### v1

# The Universal Weight Subspace Hypothesis

Abstract
We show that deep neural networks trained across diverse tasks exhibit remarkably similar low-dimensional parametric subspaces. We provide the first large-scale empirical evidence that demonstrates that neural networks systematically converge to shared spectral subspaces regardless of initialization, task, or domain. Through mode-wise spectral analysis of over 1100 models - including 500 Mistral-7B LoRAs, 500 Vision Transformers, and 50 LLaMA-8B models - we identify universal subspaces capturing majority variance in just a few principal directions. By applying spectral decomposition techniques to the weight matrices of various architectures trained on a wide range of tasks and datasets, we identify sparse, joint subspaces that are consistently exploited, within shared architectures across diverse tasks and datasets. Our findings offer new insights into the intrinsic organization of information within deep networks and raise important questions about the possibility of discovering these universal subspaces without the need for extensive data and computational resources. Furthermore, this inherent structure has significant implications for model reusability, multi-task learning, model merging, and the development of training and inference-efficient algorithms, potentially reducing the carbon footprint of large-scale neural models.
Figure 1: Deep Networks Converge to Shared, Low-Rank (Universal) Subspaces. Across distinct architectures and modalities, neural networks systematically learn to operate within remarkably similar low-dimensional parameter subspaces. Left: Principal component analysis of 200 GPT2, 500 Vision Transformers, 50 LLaMA-8B, and 8 Flan-T5 models reveals consistent sharp spectral decay - strong evidence that a small number of weight directions capture dominant variance despite vast differences in training data, objectives, and initialization. The black baseline (independent subspaces reference) represents the naive expectation that models would learn distinct directions; our empirical findings contradict this. Right: Strikingly, 500 randomly initialized ViT models converge to a common low-rank subspace, demonstrating this is a fundamental neural network property. This emergent structure unlocks powerful applications: parameter-efficient adaptation, efficient model merging, compressed storage, and accelerated training and inference. Further discussion in Figure˜7.
1 Introduction
We show that backpropagated neural networks trained on a variety of datasets - which could be disjoint and unrelated - diverse hyper-parameter settings, initializations and regularization methods, often learn an architecture-specific, layer-wise similar, low-rank joint subspaces (we refer to this as the Universal Subspace). We provide the first large-scale empirical analysis - across a diverse set of models - that neural networks tend to converge to these joint subspaces, largely independent of their initialization or the specific data used for training. Our study encompasses different model architectures trained on a variety of datasets, sometimes with different loss functions and tasks. Our spectral subspace analysis of the weights of all these models (Figure 1) suggests that although individual tasks appear to induce distinct subspaces, individually, they are all part of an unusually low-ranked joint subspace. Our work extends the scientific community’s understanding of what neural networks learn. This universality could explain several puzzling neural properties: why overparameterized models with millions more parameters than training samples still generalize; how different initializations converge to similar representations; and why techniques like weight sharing and parameter-efficient fine-tuning succeed across architectures. If networks indeed learn within shared subspaces, this would provide a supporting explanation for implicit regularization, transferability, and the effectiveness of sparse training methods, while also opening up avenues for applications like efficient merging, new optimization techniques, faster and more efficient learning and inference.
Several lines of prior research have hinted at phenomena consistent with our joint (universal) subspace hypothesis. For example, Neural Tangent Kernel (NTK) theory demonstrates that, in the infinite-width limit, the training dynamics of deep networks are governed by a kernel largely invariant to task specifics (jacot2018ntk). Similarly, research into mechanistic interpretability—specifically its own universality hypothesis (olah2020zoom; chughtai2023toymodeluniversalityreverse)—has uncovered recurring circuits and patterns within specific layers of toy or vision networks, lending indirect support to the concept of universality. Other phenomena, such as the lottery ticket hypothesis (frankle2019lottery) and mode connectivity (garipov2018mode), provide further evidence for the existence of reusable, low-dimensional structures in neural networks. notably, Krizhevsky2012ImageNet observed that the first layer of convolutional networks consistently learns Gabor-like filters across diverse vision tasks. More recently, studies by guth2024universalityneuralencodingscnns; guth24; kaushik2025eigenloraxrecyclingadaptersprincipal have shown initial evidence of recurring eigenvectors in certain layers of convolutional neural networks trained on natural images. While works such as martin2025setolsemiempiricaltheorydeep; Mao_2024 have explored neural properties to explain generalization or loss landscape convergence, they do not address the convergence of parametric properties across distinct models trained on disjoint data. Importantly, in contrast to the abstract, weak, or speculative notions of universality presented in earlier works like chughtai2023toymodeluniversalityreverse; olah2020zoom - which focus primarily on representations and are arguably easier to demonstrate due to direct data dependency - ours is the first work to provide concrete evidence and a clear universal hypothesis at the neural parameter or weights level.
In our analysis, we present compelling empirical evidence for the existence of universal subspaces within LoRA adapters across different modalities and tasks. We initially focus on LoRA adapters due to their ease of training and the ability to collect a large number of adapters for diverse tasks, models, and datasets, which enables robust evaluation of our hypothesis. E.g., we demonstrate the emergence of a universal subspace across approximately 500 LoRA adapters for the Mistral-7B (jiang2023mistral7b) model. We further extend our investigation to the full weight space, where we observe similar universality, extracting sparse, low-rank universal subspaces from about 500 Vision Transformer models and 50 LLaMA3-8B models, each trained on different datasets and initializations.
Although the underlying causes and broader implications of this universal property remain an open area of investigation, even an initial understanding of parameter subspace universality has profound implications for neural network efficiency and interpretability. Shared subspaces could enable: (1) massive model compression by storing only subspace coefficients rather than full weights; (2) rapid adaptation to new tasks within learned subspaces; (3) theoretical insights into generalization bounds and optimization landscapes; and (4) environmental benefits through reduced computational requirements for training and inference.
To date, our work presents the most rigorous empirical evidence for the existence of “Universality” within the parameter space of deep neural networks. This geometric universality offers a novel vantage point for investigating fundamental neural properties, including generalization, grokking, catastrophic forgetting, and dataset efficiency. Crucially, the pervasive nature of this universal property underscores the relative primacy of model architecture over other factors in shaping the learned parameter space (.) However, our findings also delineate clear frontiers for future inquiry. We leave open the question of cross-architectural comparison: how do the universal subspaces of distinct architectures differ, and can we explicitly design architectures to optimize the geometry of this subspace? Furthermore, a fundamental question emerges from our results regarding the implications of convergence: if neural networks systematically collapse into the same subspace - thereby inheriting shared biases, capabilities, and failure modes - is this lack of diversity a fundamental bottleneck, and should we develop methods specifically designed to break this convergence?
The remainder of this paper is organized as follows. We first define the problem set up formally in Section 2 followed by listing of essential properties and conditions with corresponding empirical justifications. Section 3.3.1 proposes the method to adapt to new tasks leveraging the shared approximate universal subspace. Section 3.1 explains our analysis methodology and Section 3.2 presents the comprehensive empirical evidence of the Universal subspaces. Section 4 briefly discusses the analysis providing useful insights and answers the fundamental questions raised in the introduction. We discuss related work in appendix A.1 and discuss limitations and scope for future work in Section 5. Our primary contributions include
  • We empirically demonstrate the existence of a lower-dimensional shared universal subspace in backpropagated neural networks, and also provide relevant theoretical analysis.
  • Illustrate the approach to learning an approximate low-dimensional shared subspace using the available set of tasks. Propose conditions for convergence of this learned subspace to the true universal shared subspace.
  • Reuse the learned shared subspace to efficiently adapt to new unseen tasks with significantly fewer of trainable parameters. Our experiments across a wide variety of large pretrained models across various architectures and data modalities extensively verify and validate our hypothesis and theoretical findings.
  • We show that we can use the Universal subspace for efficient, faster learning, efficient model scaling, and model compression.
2 Notations, Definitions and Theoretical Analysis
Our theoretical analysis models predictors as elements of a Hilbert space, for example a reproducing kernel Hilbert space (RKHS), while our experiments are conducted with practical large-scale models such as transformers and LoRA-based variants. Modeling predictors in a Hilbert space (kernel) framework is standard when analyzing aspects such as generalization and inductive bias of modern deep architectures, and has been widely used to approximate or interpret the behavior of large neural networks in practice (ortiz-jimenez2023taskarithmetic; wei2019regularization; chen2021deep; belfer2024spectral; bietti2019kernelperspectiveregularizingdeep). We aim to understand whether the shared structure across tasks can be consistently recovered from data as number of tasks increase. Specifically, each task has an associated ground-truth predictor $f_{t}^{\star}$, and we are interested in the covariance (second-moment) operator $\mathcal{S}$ that captures the common subspace spanned by these predictors. Since in practice we only observe finite samples per task and learn approximate predictors $\hat{f}_{t}$, two sources of error arise: (i) variability due to having finitely many tasks, and (ii) estimation noise within each task. Our goal is to establish conditions under which the empirical operators built from $\hat{f}_{t}$ concentrate around $\mathcal{S}$, and to show that the learned top-$k$ subspace converges to the true one, with convergence rates that separately reflect the number of tasks and the accuracy of per-task learning.
Setup.
Let $(\mathcal{H},\left\langle\cdot,\cdot\right\rangle)$ be a separable Hilbert space with norm $\left\lVert\cdot\right\rVert=\left\lVert\cdot\right\rVert$$_{\mathcal{H}}$. For $a,b\in\mathcal{H}$, the rank-one operator $a\otimes b:\mathcal{H}\to\mathcal{H}$ is $(a\otimes b)g=\left\langle b,g\right\rangle\,a$; in particular $\left\lVert a\otimes b\right\rVert_{\mathrm{op}}=\left\lVert a\right\rVert\,\left\lVert b\right\rVert$. Tasks $t=\{1,2,3...,T\}$ are drawn i.i.d. from distribution $\mathcal{T}$  and each task dataset $S_{t}=\{(x_{t,i},y_{t,i})\}_{i=1}^{n_{t}}$ with $n_{t}$ samples is drawn independently from $D_{t}$. Let $f_{t}^{\star}\in\mathcal{H}$ denote the (unknown) ground-truth predictor for task $t$ and $\hat{f}_{t}\in\mathcal{H}$ be the learned predictor for the task.
Definition 2.1 (Task second-moment operator).
The population, true empirical, and learned empirical task second-moment operators are respectively,
$\mathcal{S}:=\mathbb{E}_{t\sim\tau}[\,f_{t}^{\star}\otimes f_{t}^{\star}\,],\qquad\hat{\mathcal{S}}:=\frac{1}{T}\sum_{t=1}^{T}f_{t}^{\star}\otimes f_{t}^{\star},\qquad\tilde{\mathcal{S}}:=\frac{1}{T}\sum_{t=1}^{T}\hat{f}_{t}\otimes\hat{f}_{t}.$
[…]
Remark 2.2.
We work with the second-moment operator (rather than centered covariance), so the top eigenspace may include the mean direction of $\{f^{\star}_{t}\}_{t\sim\mathcal{T}}$.
Let $\lambda_{1}\geq\lambda_{2}\geq\cdots$ be the eigenvalues of $\mathcal{S}$ with orthonormal eigenvectors $\{\phi_{i}\}_{i\geq 1}$. Write $P_{k}=\sum_{i=1}^{k}\phi_{i}\otimes\phi_{i}$ for the projector onto the population top-$k$ subspace $\mathcal{H}_{k}^{\star}=\mathrm{span}\{\phi_{1},\dots,\phi_{k}\}$, and let $\tilde{P}_{k}$ be the projector onto the top-$k$ eigenspace of $\tilde{S}$ (the learned shared subspace). Define the eigengap $\gamma_{k}:=\lambda_{k}-\lambda_{k+1}>0$.
Assumption 2.3 (Realizability, bounded second moment and effective rank).
For a constant $B>0$ and for all tasks, $f_{t}^{\star}\in\mathcal{H}$ almost surely, $\left\lVert f_{t}^{\star}\right\rVert\leq B$ a.s., $\mathbb{E}_{t\sim\tau}\left\lVert f_{t}^{\star}\right\rVert^{2}=\operatorname{tr}(S)<\infty$. In addition, $\mathcal{S}$ has bounded effective rank, $\frac{tr(\mathcal{S})}{\left\lVert\mathcal{S}\right\rVert_{\mathrm{op}}}\leq\kappa$
Assumption 2.3 ensures that all ground-truth predictors are bounded and have finite second moment, so the population covariance operator $S$ is well-defined. The bounded effective rank condition further guarantees that the shared structure of the tasks is not arbitrarily infinite-dimensional, making subspace recovery feasible.
Assumption 2.4 (Per-task estimation accuracy in $\mathcal{H}$).
For any $\delta_{t}\in(0,1)$ with probability at least $1-\delta_{t}$ over the draw of $S_{t}$,
$\left\lVert\hat{f}_{t}-f_{t}^{\star}\right\rVert\ \leq\ \eta_{t},\text{\hskip 5.69054pt...where }\eta_{t}=\mathcal{R}_{n_{t},D_{t}}(\mathcal{H})+\sqrt{\frac{\ln(1/\delta_{t})}{2n_{t}}}$
Here $\mathcal{R}_{n_{t},D_{t}}(\mathcal{H})$ represents Rademacher complexity of the solutions within Hilbert space $\mathcal{H}$ over $n_{t}$ samples drawn i.i.d. from $D_{t}$ This form is satisfied, for example, by strongly convex regularized ERM in an RKHS (e.g., kernel ridge regression or NTK ridge), under bounded kernel norm and sub-Gaussian response noise (barlette2003rademacher).
Assumption 2.4 requires that each task predictor $\hat{f}_{t}$ is learned accurately from its finite dataset. In other words, $\hat{f}_{t}$ is close to the true $f_{t}^{\star}$ in $\mathcal{H}$-norm with high probability, at a rate governed by sample size and complexity of the hypothesis space.
Theorem 2.5 (Two-level convergence to the shared subspace).
Assume 2.32.4. Let $c_{1},c_{2}$ be any absolute constants. For any $\delta\in(0,1)$, choose $\delta_{t}=\delta/(2T)$ and set $\delta_{T}=\delta/2$. With probability at least $1-\delta$ (over tasks and all per-task samples),
$\left\lVert\tilde{\mathcal{S}}-\mathcal{S}\right\rVert_{\mathrm{op}}\ \leq\ c_{1}B^{2}\sqrt{\frac{\log(c_{2}/\delta)}{T}}\ +\ (2B\bar{\eta}+\overline{\eta^{2}})$ (1)
If moreover $\gamma_{k}>0$, then
$\left\lVert\tilde{P}_{k}-P_{k}\right\rVert_{\mathrm{op}}\ \leq\ \frac{2}{\gamma_{k}}\!\left(c_{1}B^{2}\sqrt{\frac{\log(c_{2}/\delta)}{T}}+(2B\bar{\eta}+\overline{\eta^{2}})\right).$ (2)
where $\bar{\eta}=\frac{1}{T}\sum^{T}_{t=1}\eta_{t}$, $\overline{\eta^{2}_{t}}=\frac{1}{T}\sum^{T}_{t=1}\eta^{2}_{t}$ and $\eta_{t}$ is defined same as in assumption 2.4
Proof of Theorem˜2.5 can be found in appendix Section˜A.2. The Theorem˜2.5 shows that the empirical second-moment operator built from the learned predictors converges to the true operator $\mathcal{S}$, and the learned top-$k$ subspace $\hat{P}_{k}$ converges to the true subspace $P_{k}$. The rates capture two sources of error: averaging across tasks (scaling with $1/\sqrt{T}$) and per-task estimation errors (through $\bar{\eta}$ and $\overline{\eta^{2}}$). A larger eigengap $\gamma_{k}$ makes the subspace recovery more stable. In practice, we obtain the eigenvectors of $\tilde{\mathcal{S}}$ using HOSVD (Higher-Order Singular Value Decomposition) of the concatenated weight matrix $\mathcal{X}$ highlighted in Section˜3. Motivated by our theoretical analysis, we try to approximate $\hat{\mathcal{S}}$ for a set of tasks by extracting principal directions from as many trained models as possible.
3 Analysis
This analysis constitutes the core contribution of our work. We demonstrate that architecture-specific, layer-wise universal subspaces consistently emerge across a diverse array of neural models. In our experiments, adherence to the universal subspace hypothesis appears robust, showing no significant deviation regardless of whether models are trained from scratch, fully finetuned, or adapted via low-rank methods. Furthermore, this phenomenon remains invariant across differing initialization strategies, modalities, data formats, and dataset contents. Notably, however, the fidelity of the extracted subspace correlates with the quantity and quality of the available models. This observation leads us to postulate the existence of an "ideal" universal subspace intrinsic to each architecture, towards which individual model instances converge. We hypothesize that superior algorithms, cleaner data, and more effective optimization strategies enable models to approximate this ideal more closely. While we do not formally verify the "ideal universal subspace" hypothesis in the present work, leaving it for future investigation, we suggest that this subspace represents the most stable configuration for contemporary neural networks trained via backpropagation. Consequently, exceptions to this rule may offer fertile ground for further research. We present the analysis methodology in Section˜3.1, and experimental evidence of the universal subspace using the methodology in the following subsections.
3.1 Analysis methodology
Algorithm 1 Truncated Zero-Centered Higher-Order SVD (HOSVD)
1:A high-order tensor $\mathcal{X}\in\mathbb{R}^{I_{1}\times\cdots\times I_{N}}$ constructed by stacking NN rank-$r_{n}$ task matrices along mode $n$, where $1\leq r_{n}\leq I_{n}$ and $n\in[1,N]$.
2:Mean tensor $\bm{\mu}$; factor matrices $U^{(n)}\in\mathbb{R}^{I_{n}\times\hat{r}_{n}}$ (orthonormal columns), where $\hat{r}_{n}$ is chosen as the smallest number of left singular vectors whose cumulative explained variance is at least $\tau$; and the truncated core tensor $\mathcal{S}\in\mathbb{R}^{\hat{r}_{1}\times\cdots\times\hat{r}_{N}}$. Reconstruction is given by $\widehat{\mathcal{X}}=\bm{\mu}+\mathcal{S}\mathbin{\times_{1}}U^{(1)}\cdots\mathbin{\times_{N}}U^{(N)}$, where $\mathbin{\times_{n}}$ denotes mode-$n$ tensor–matrix multiplication.
3:Zero-centering: $\bm{\mu}\leftarrow\mathrm{mean}(\mathcal{X})$ $\triangleright$ elementwise mean over all entries
[…]
5:for $n=1$ to NN do
6:  $X_{(n)}\leftarrow\mathrm{unfold}(\mathcal{X}_{c},\,n)$ $\triangleright$ mode-$n$ matricization; $X_{(n)}\in\mathbb{R}^{I_{n}\times\prod_{m\neq n}I_{m}}$
7:  Compute thin SVD: $X_{(n)}=\tilde{U}^{(n)}\Sigma^{(n)}\tilde{V}^{(n)\top}$
8:  $U^{(n)}\leftarrow\tilde{U}^{(n)}(:,\,1{:}\hat{r}_{n})$ $\triangleright$ keep first $\hat{r}_{n}$ left singular vectors (variance $\geq\tau$)
9:end for
10:Truncated core: $\mathcal{S}\leftarrow\mathcal{X}_{c}\mathbin{\times_{1}}U^{(1)\top}\mathbin{\times_{2}}U^{(2)\top}\cdots\mathbin{\times_{N}}U^{(N)\top}$
11:return $\bm{\mu},\,\{U^{(n)}\}_{n=1}^{N},\,\mathcal{S}$ $\triangleright$ Optionally compute $\widehat{\mathcal{X}}=\bm{\mu}+\mathcal{S}\mathbin{\times_{1}}U^{(1)}\cdots\mathbin{\times_{N}}U^{(N)}$
Since there is no current method that enables us to compare subspaces of models with different architectures, we focus on large number of models trained on the same architecture. To this end, we perform analysis using Low rank adapters (hu2021lora) (LoRA) as well as classical weights of transformer and CNN (Convolutional Neural Network) architectures. For all our experiments, unless stated otherwise, we perform Order 1-2 HOSVD only, to ensure that our methodology works even in the simplest case. Algorithm˜1 provides the algorithm we implement. Refer to Appendix˜B for discussion regarding secondary subspace and how to choose the number of top components.
3.2 Results From Joint Subspaces’ Analysis
We present empirical results using method shown in Section 3.1, extracting our layer wise universal subspace approximations using thousands of publicly available models for most of our experiments. This choice allows us to have no training costs whatsoever, for extracting the universal subspace. Spectral analysis relies on efficient spectral decomposition libraries, and can even be run on CPUs. We run all our analysis and experiments on one Nvidia A5000 GPU. The presented large scale empirical results forms the crux of our work and provide strong evidence for the presence of such low ranked joint subspaces across a wide range of task, architecture and modalities. In summary, we present a total of eight set of analysis and applications, including tasks like image classification, natural language understanding, text to image generation, model merging, etc for different model architectures and modalities.
3.2.1 Lower-rank joint subspaces in CNNs, LoRA and Finetuned models
In smaller and conventional architectures such as CNNs, evidence for universal structure has been more limited but suggestive. Early work observed that the first convolutional layer often learns Gabor-like filters across diverse vision tasks (Krizhevsky2012ImageNet). More recently, works report recurring eigenvectors in certain CNN layers trained on natural images (guth24; guth2024universalityneuralencodingscnns).
We extend these observations and examine whether a shared low-rank joint subspace emerges across tasks. Specifically, we train ResNet-50 models from random initialization for image classification on five disjoint datasets (CIFAR-10, CIFAR-100, ImageNet, Oxford-IIIT Pets, and EuroSAT), ensuring no overlap in samples. While our theoretical analysis indicates that a small number of models may lead to an under-approximation of the joint universal subspace, training CNNs from scratch at scale constrains the number of models we can include in this study.
(a) Comparison of model performance across datasets.
Method ImageNet EuroSat CIFAR-10 CIFAR-100 Oxford Pets Avg
ResNet50 80.86 98.96 97.35 83.82 93.48 90.89
Universal R50 77.89 98.83 95.89 81.49 83.81 87.58
(b) Summarized (averaged for all layers) eigenvalue plot of all model weights corresponding to all 31 layers of 5 ResNet50 models. Mean refers to the fact that it has been averaged for all layers for conciseness. The vertical axis is Explained Variance (for all models) and X axis indicated Principal Components. We will follow this setup throughout the paper. We also refer to the low-ranked shared subspace as ’Universal’ subspace and may refer to a specific model consisting of extracted basis as the ’Universal variant’.
Figure 2: Proving existence of universal subspaces in CNNs. Decomposing 5 ResNet50 models trained on different tasks shows the emergence of a low rank, universal subspace where the majority of the information is present in only 16 (or fewer) distinct subspace directions for all layers of the network.
Despite these limitations, Figure 2(b) reports the average explained variance across all layers of ResNet-50 and reveals a distinct, shared low-rank structure spanning these disjoint tasks. Moreover, even when the estimated universal subspace is relatively coarse, projecting to this subspace to obtain a low-rank ResNet-50 (thereby reducing parameters) preserves competitive performance relative to full fine-tuning, further supporting the presence and utility of a joint subspace (2(a)).
In order to conduct a more real-world experiment, we choose to run the subspace analysis for LoRA hu2021lora models simply because they are available in abundance in public domain. Given LoRA models distinctly capture task specific directions as they show weak alignment with the original weights hu2021lora, they form a good main model parameter alternative to run our subspace analysis and verify whether this holds true. We spectrally decompose (Section 3.1) LoRA’s submatrices individually, each concatenated across all the available finetuned LoRAs and choose top $k$ spectral basis. This setup allows us to truly stress test the Universal Subspace.
(a) Eigenvalue/Variance plot for Orthogonal Spectral Components for 500 unique LoRAs of different layers of Mistral-7B model
(b) Summarized eigenvalue plot of all LoRAs corresponding to all 31 layers of all 500 Mistral 7B models
Figure 3: Proving existence of universal subspaces in deep networks. Decomposing 500 sets of LoRAs trained on different tasks using the Mistral-7B model shows the emergence of a low rank, universal subspace where the majority of the information is present in only 16 (or less) distinct subspace directions for all layers of the network. Plots of other layers are present in the Section˜B.1.
In our first experiment, we use LoRA models trained on 500 natural instruction tasks (wang-etal-2022-super) using Mistral-7B-Instruct-v0.2 (jiang2023mistral7b) as the base (brüelgabrielsson2024compressserveservingthousands). Each LoRA, individually, is at least of rank 16. Figure 3 shows the results of our analysis. 3(a) shows the subspace analysis of individual layers of the Mistral model. Each bar corresponds to the Eigenvalue or explained variance of every unique component (unique for all models, trained in disjoint/different datasets), showing that the parameters of all 500 models can be well approximated by a finite low-rank subspace for all layers. Note that for visualization purposes, the graphs only show the top 100 components, but there exists a significantly larger number of basis. 3(b) provides a summary of this subspace analysis, (summarized) for all of the layers of the 500 Mistral models, further validating the idea of a layer-wise universal subspace. Please refer to the appendix for all other and larger versions of the given spectral plots.
To test the expressiveness of this universal subspace, we analytically reconstruct the LoRA parameters of randomly chosen seen (IID) and unseen (OOD) tasks by projecting it onto the universal subspace. Figure 4 shows the result of our experiment, and as can be seen, that the Universal subspace model perform robustly for both seen and unseen cases. In order to verify the importance of this chosen subspace, we redo the experiment with the leftover components from our spectral decomposition (called Secondary Subspace). As can be seen from the results, the performance of this model lags drastically behind our chosen subspace. We note that our universal subspace model is $19\times$ more memory efficient since it is no longer necessary to save all 500 LoRAs.
Figure 4: Lots of LoRAs Model Size vs Performance plot.
We extend our analysis to text-to-image generation using Stable Diffusion-XL (sdxl). A universal subspace is extracted from publicly available LoRAs on HuggingFace (von-platen-etal-2022-diffusers). When projecting individual LoRAs into this subspace, the resulting generations preserve visual quality and style (Figure 5). CLIP-based evaluations (Table 1) show that the universal subspace even outperforms individual LoRAs in some cases, possibly due to denoising effects previously observed in (sharma_laser_2023).
Table 1: CLIP scores (higher is better) of images generated using SDXL.
Method Style 1 Style 2 Style 3 Style 4 Style 5 Style 6 Style 7 Style 8 Style 9 Style 10 Avg
LoRA 21.95 15.59 22.18 18.84 16.65 17.99 24.66 17.47 22.07 19.93 19.73
Universal SDXL LoRA 21.96 16.07 22.07 18.79 16.68 17.99 24.66 17.56 22.46 20.09 19.83
Figure 5: Text-to-Image Generation Results for Individual models vs. our Universal Subspace model. We notice no visual reduction in style quality despite significant reduction in total model size.
In order to test the ability of condensing many models into a single universal subspace, we compare our method with SOTA model merging/combination methods in Table 2. We compare our universal subspace inspired combination approach against six state-of-the-art, gradient-free baselines: RegMean (jin2023dataless), Task Arithmetic (TA) (ilharco2023editing), TIES (yadav2023tiesmerging), DARE-TIES (YuDare), KnOTS-TIES, and KnOTS-DARE-TIES (knots). RegMean aligns task-specific updates by solving a layer-wise linear regression problem, requiring transformation matrices for each model. TA merges models by linearly combining parameters, but relies on tuning scaling coefficients on a validation set for optimal performance. TIES extends TA with magnitude-based pruning and sign conflict resolution, introducing additional hyperparameters such as pruning thresholds, while DARE-TIES combines random Bernoulli pruning with TIES’ sign resolution, also requiring tuning of pruning probability. KnOTS-TIES and KnOTS-DARE-TIES further apply SVD-based subspace alignment before merging, but still inherit the need for coefficient or pruning hyperparameter selection. In contrast, our universal subspace method, analytically computes the merging coefficients based solely on the geometry of a shared, low-rank universal subspace identified across models, requiring no iterative tuning or validation data-although optional finetuning is possible if data is available. Furthermore, because our subspace is intrinsically low-rank, the merged model contains significantly fewer parameters than any individual models, offering both computational efficiency and theoretical alignment guarantees not present in the baselines. Empirically, our approach achieves higher average accuracy (see Table 2), while reducing parameter count, thus enabling scalable and robust model merging without heuristic pruning or validation overhead. We note that we did not optimize our merging process and better results nearing finetuned performance may be achieved.
Table 2: Per-task results for eight ViT-B/32 models, each finetuned with LoRA on a different image classification dataset. "Finetuned" indicates the accuracy of each model on its respective training dataset. For each merging baseline, we report the normalized accuracy on every task, as well as the average across all tasks.
Method Datasets Avg
Cars DTD EuroSAT GTSRB MNIST RESISC45 SUN397 SVHN
Per-Task Absolute Accuracies (%)
Finetuned 74.0 58.3 99.0 92.7 99.3 88.4 64.5 96.2 84.1
Per-Task Accuracies of Combined Models Normalized Against Finetuned Models (%)
RegMean 80.2 71.3 37.9 47.3 43.1 70.5 99.3 43.0 60.9
TA 82.0 73.6 48.8 42.1 53.1 71.5 97.5 41.2 63.7
TIES 82.4 72.8 50.8 39.0 50.3 70.9 99.4 40.5 63.7
DARE-TIES 81.4 74.5 50.8 39.2 55.0 70.7 96.7 40.4 63.7
KnOTS-TIES 82.7 73.7 49.3 48.9 70.9 95.5 53.8 68.0 68.0
KnOTS-DARE-TIES 81.8 75.9 50.7 40.3 53.2 70.2 97.9 41.0 63.9
Ours 88.1 82.3 65.9 61.3 88.3 98.1 98.5 85.1 83.5
In summary, these four experiments provide strong empirical support for our universal subspace hypothesis and demonstrate its practical advantages in terms of memory efficiency, model merging, model reusability, and scalable deployment across diverse tasks and modalities.
3.3 Low rank shared universal subspaces in classical weights
Table 3: Image Classification Accuracy
Method IID OOD
Full Training 94.4 $\pm$ 1.7 91.3 $\pm$ 2.1
Universal ViT 94.1 $\pm$ 2.0 87.8 $\pm$ 1.5
While aforementioned experiments on CNNs trained from scratch, and LoRAs, provide strong evidence for the presence of the joint subspace, we further rigorously test on large scale finetuned models (500 pretrained ViT, 50 LLaMA3-8B models, 177 GPT-2 and Flan-T5).
Figure 6: Universal Subspaces in Classical Weights. Spectral decomposition of weight matrices from (a) $\sim$500 Vision Transformers (b) 50 LLaMa-8B models (c) 177 GPT-2 models (d) GLUE Flan-T5 models - each trained independently across diverse tasks, datasets, and configurations - reveals a consistent low-rank structure: most variance is captured by the top few spectral basis. This suggests that, despite significant variation in training conditions, the learned weights consistently align along a shared low-dimensional subspace. For visualization clarity, only a fraction of the basis are shown; extended plots are provided in the Section˜B.2.
First, we collect $\sim$500 pretrained Vision Transformer (ViT) models from HuggingFace, spanning diverse domains - medical imaging, satellite data, and synthetic - and trained with varying losses, optimizers, and initializations. Importantly, we did not filter or curate these models in any way and had no access to the original training data, ensuring our analysis reflects real-world diversity. Details about the models and their configurations are provided in Section˜B.2. Following our method (Section˜3.1), we spectrally decompose all layers (excluding first and last) and observe, in Figure 6, that the majority of variance is captured by the top few spectral components, revealing a highly compressible, shared subspace across layers. Only the top 100 components are visualized for clarity.
To further validate the universality of this subspace, we selected five additional, previously unseen pretrained ViT models for which we had access to evaluation data. These models, considered out-of-domain relative to the initial set, had all their weights reconstructed by projecting onto the identified 16-dimensional universal subspace. We then assessed their classification accuracy and found no significant drop in performance, as reported in Table 3. This result provides strong empirical support for our universal subspace hypothesis, demonstrating that a shared, low-dimensional structure underlies the weight spaces of diverse, independently trained Vision Transformers, regardless of their training data or configuration.
[…]
We further extend this analysis to 50 finetuned LLaMA3-8B models, 177 GPT-2 models, and Flan-T5 models (trained on GLUE glue datasets) again sourced from HuggingFace without filtering. As shown in Figure 6, a small number of directions capture dominant structure across models spanning diverse and distinct datasets and tasks. More details are provided in the Section˜B.2. This is, to our knowledge, the first instance of compressing such a large and diverse collection of foundation models into a unified subspace, highlighting its potential for large-scale model reuse and environmental efficiency.
3.3.1 Finding universal subspaces and applying them to future tasks
In this section, the low-rank shared subspaces estimated from a set of available tasks are leveraged to adapt to new, previously unseen tasks. While we do not make theoretical guarantees about reuse on unseen tasks, our experiments show that the approximate shared subspace is empirically reusable across a wide range of practical settings. Concretely, we reuse the shared principal directions and learn only their task-specific coefficients for the new task. Learning these low-rank coefficients is substantially cheaper than optimizing full-rank weights of size, reducing both computation and memory. The resulting trainable parameter counts are reported in Table 5. We find our universal subspace models can have significant impact on the carbon footprint issues of large AI models by making the training, inference and scaling of these models efficient and cheap. As shown in the previous section, we can effectively recycle and replace available pretrained models with a universal subspace model with every individual being represented by a sparse set of coefficients. In this section, we show a set of experiments where we utilize the universal subspaces to learn new tasks by freezing the components and simply learning the coefficients using gradient descent. We find that since we are only learning the coefficients, it drastically cuts down the number of parameters required to train the new models. Further, since these coefficients are simply linear scaling values, the optimization is smoother and faster.
Table 4: Performance on the GLUE Benchmark.
Method Speedup CoLA MRPC RTE QNLI SST-2 STS-B Avg
LoRA $1\times$ 59.56 86.76 77.61 92.53 94.72 90.81 83.67
Universal order-2 $2\times$ 61.82 87.25 77.62 92.71 94.15 90.48 84.01
Universal order-3 $1.8\times$ 62.06 86.52 75.81 92.98 94.26 90.39 83.67
We present two experiments - Image Classification using ViT-base and Natural Language Understanding using GLUE benchmark glue with RoBERTa-base model. Both involve creating a universal subspace using publicly available LoRA adapters. Details are provided in the Appendix˜C. For the GLUE benchmark, we follow the same setup as  (kopiczko_vera_2023) considering the 6 tasks - CoLA, MRPC, SST-2, QNLI, RTE and STS-B while omitting the time-intensive MNLI and QQP tasks. We initialize our universal subspace using a leave-one-out-setup, where the subspace is calculated using components of all but one LoRA adapter for which the coefficients are learned. For image classification, we utilize publicly available ViT LoRAs to extract our universal subspaces taking care that the data any of these pretrained LoRAs have not seen the data we will be training our coefficients on.
Table 5: Image Classification with Vision Transformer.
# Training Params CIFAR100 Food101 Flowers102 CIFAR10 Pets
Full Training 86M 92.8 90.7 98.82 99.0 91.2
Universal ViT 10K 90.1 89.1 90.1 96.7 89.4
[…]
4 Discussion
This work provides, to the best of our knowledge, the first large-scale, cross-domain analysis showing that neural networks trained across diverse tasks, modalities, initializations, and hyperparameters consistently exhibit an architecture-specific shared low-rank universal subspace at the layer level. Concretely, by performing layer-wise spectral decompositions and retaining only the leading principal directions, an accurate approximation of these universal subspaces can be extracted. Empirically, this behavior emerges broadly: in fully finetuned models and LoRA-based adapters, in models trained from scratch, in both generative and discriminative settings, and in multimodal configurations. Moreover, the approximated subspaces generalize to out-of-distribution tasks, where projecting models and learning only a small set of coefficients suffices to recover strong performance. This enables adapting to new tasks without retraining or storing full weights, and supports robust multi-task learning, scalable fine-tuning, and principled model merging within a single unifying framework.
The practical implications are substantial. By reusing a common set of layer-wise principal directions and learning only lightweight coefficients per task, large models can be extended and served with dramatically reduced computational, memory, and engineering overhead. This directly lowers both the financial and environmental costs of training and deployment, aligning with the broader goals of sustainable and accessible AI. Reducing hardware and energy requirements for adaptation and inference opens participation to under-resourced researchers and institutions while facilitating modular design, data-free or data-minimal model merging, and more maintainable systems. Taken together, these results suggest a path toward scalable, equitable, and interpretable model reuse grounded in a simple geometric principle: most task variation lies in a shared, low-dimensional subspace.
Why do these universal subspaces emerge? While the precise mechanisms driving this phenomenon remain an open area of investigation, several theoretical factors likely contribute to the emergence of these shared structures. First, neural networks are known to exhibit a spectral bias toward low-frequency functions, creating a polynomial decay in eigenvalues that concentrates learning dynamics into a small number of dominant directions (belfer2024spectral; bietti2019kernelperspectiveregularizingdeep). Second, modern architectures impose strong inductive biases that constrain the solution space: convolutional structures inherently favor local, Gabor-like patterns (Krizhevsky2012ImageNet; guth24), while attention mechanisms prioritize recurring relational circuits (olah2020zoom; chughtai2023toymodeluniversalityreverse). Third, the ubiquity of gradient-based optimization – governed by kernels that are largely invariant to task specifics in the infinite-width limit (jacot2018ntk) – inherently prefers smooth solutions, channeling diverse learning trajectories toward shared geometric manifolds (garipov2018mode). If these hypotheses hold, the universal subspace likely captures fundamental computational patterns that transcend specific tasks, potentially explaining the efficacy of transfer learning and why diverse problems often benefit from similar architectural modifications.
5 Limitations and Future Work
Although we provide conclusive results towards the existence and utility of universal shared subspaces, the current analysis has scope for future research, such as limited interpretability of the shared subspace and the corresponding directions. While it is a critical area of research, it is extremely cumbersome to demonstrate interpretability of the principal directions for each layer of the network. To the best of our knowledge we are not aware of any other literature that performs such an in-depth analysis of the weight space of large models across diverse tasks, data modalities and model architectures. The current approach to approximating a universal subspace relies on pretrained task-specific models (predictors) for tasks, which may not be readily available for new tasks. An interesting direction for future research would be to explore model independent methods for learning a universal shared subspace, potentially derived directly from data. Furthermore, the conditions proposed in ortiz-jimenez2023taskarithmetic for enabling task arithmetic rely on localized eigenfunctions which are not conducive to learning a shared universal subspace. As a result, performing task arithmetic within the current framework of a shared universal subspace is non-trivial and warrants further investigation.
Appendix A Appendix
Table 6: Notation reference.
Notation Description
$\mathcal{H}$ Separable Hilbert space with inner product $\langle\cdot,\cdot\rangle$, norm $\|\cdot\|$.
$a\otimes b$ Rank-one operator $g\mapsto\langle b,g\rangle a$, $\|a\otimes b\|_{\mathrm{op}}=\|a\|\,\|b\|$.
$T$ Number of tasks.
$\mathcal{T}$ Distribution over tasks.
$D_{t}$ Data distribution for task $t$.
$S_{t}=\{(x_{t,i},y_{t,i})\}_{i=1}^{n_{t}}$ Dataset of size $n_{t}$ for task $t$.
$f_{t}^{\star}\in\mathcal{H}$ Ground-truth predictor for task $t$.
$\hat{f}_{t}\in\mathcal{H}$ Learned predictor for task $t$.
$B$ Uniform bound: $\|f_{t}^{\star}\|\leq B$ almost surely.
$\mathcal{R}_{n_{t},D_{t}}(\mathcal{H})$ Per-task estimation error rate (e.g. $\tilde{O}(1/\sqrt{n_{t}})$).
$\eta_{t}$ Per-task error: $\eta_{t}:=\mathcal{R}_{n_{t},D_{t}}(\mathcal{H})+\sqrt{\tfrac{\ln(2T/\delta)}{2n_{t}}}$.
$\bar{\eta}$ Average error: $\tfrac{1}{T}\sum_{t=1}^{T}\eta_{t}$.
$\overline{\eta_{t}^{\,2}}$ Average squared error: $\tfrac{1}{T}\sum_{t=1}^{T}\eta_{t}^{2}$.
$\mathcal{S}$ Population operator: $\mathcal{S}=\mathbb{E}_{t\sim\mathcal{T}}[f_{t}^{\star}\otimes f_{t}^{\star}]$.
$\hat{\mathcal{S}}$ Empirical operator (true predictors): $\tfrac{1}{T}\sum_{t=1}^{T}f_{t}^{\star}\otimes f_{t}^{\star}$.
$\tilde{\mathcal{S}}$ Empirical operator (learned predictors): $\tfrac{1}{T}\sum_{t=1}^{T}\hat{f}_{t}\otimes\hat{f}_{t}$.
$\lambda_{1}\geq\lambda_{2}\geq\dots$ Eigenvalues of $\mathcal{S}$.
$\phi_{i}$ Orthonormal eigenvectors of $\mathcal{S}$.
$P_{k}$ Projector onto top-$k$ eigenspace of $\mathcal{S}$.
$\tilde{P}_{k}$ Projector onto top-$k$ eigenspace of $\tilde{\mathcal{S}}$.
$\gamma_{k}$ Eigengap: $\gamma_{k}:=\lambda_{k}-\lambda_{k+1}>0$.
$\|A\|_{\mathrm{op}}$ Operator (spectral) norm.
$\|A\|_{HS}$ Hilbert–Schmidt norm.
$r(V)$ Intrinsic/Effective rank: $\mathrm{tr}(V)/\|V\|_{\mathrm{op}}$.
$X_{t}$ Centered operator: $X_{t}:=f_{t}^{\star}\otimes f_{t}^{\star}-\mathcal{S}$.
$V$ Variance operator: $V:=\sum_{t=1}^{T}\mathbb{E}[X_{t}^{2}]$.
$\delta,\delta_{t},\delta_{T}$ Failure probabilities (global, per-task, across-task).
Figure 7: Empirical Evidence for (Universal) Joint Weight Subspaces. This figure illustrates the existence of joint low-dimensional subspaces across models trained on diverse tasks. We plot the average explained variance of the top few principal components of weight matrices from 500 Mistral-7B LoRAs, 500 Vision Transformers, and 50 LLaMA-8B models. Despite differences in modality, data, and training objective, all models exhibit rapid spectral decay - indicating that a small number of directions dominate across layers and settings. This consistent structure provides strong evidence for the presence of joint/universal subspaces, supporting our hypothesis that deep networks systematically reuse a common representational basis. Often, this shared subspace can be seen distinctly. The presence of the subspace has significant implications for deep learning. Not only can large number of models be compressed into a single, lighter Universal model with difference represented as lightweight coefficients, training on future tasks simply becomes tuning those coefficients. Since the basis are fixed, training becomes simpler and quicker. However, this convergence to similar subspace raises few important questions - is it possible to recover the "true" Universal Subspace without learning with huge amounts of data? Is this lack of diversity a bottleneck from current family of deep models?
A.1 Related Work
Several lines of prior research support the core intuition behind our universal subspace hypothesis, though they do not provide a unified, scalable framework for identifying and leveraging such subspaces across architectures, tasks, and modalities. The Neural Tangent Kernel framework reinforces this idea, demonstrating that, in the infinite-width regime, training dynamics are governed by a kernel largely invariant to task specifics, implying the presence of common functional subspaces.  (jacot2018ntk). This result implies that training is implicitly constrained to a shared function space, suggesting the existence of low-dimensional structures that generalize across tasks. Complementing this, works in mechanistic interpretability has uncovered modular and recurring patterns that consistently re-emerge in independently trained models (olah2020zoom; chughtai2023toymodeluniversalityreverse), supporting the notion of structural universality in network representations. The closest related work is our previous work kaushik2025eigenloraxrecyclingadaptersprincipal, which investigates principal subspaces in low rank adapters for highly parameter efficient finetuning. That work considered a derivation of the analysis we present in this work. While works such as martin2025setolsemiempiricaltheorydeep; Mao_2024 have explored neural properties to explain generalization or loss landscape convergence, they do not address the convergence of parametric properties across distinct models trained on disjoint data. Importantly, in contrast to the abstract, weak, or speculative notions of universality presented in earlier works like chughtai2023toymodeluniversalityreverse; olah2020zoom - which focus primarily on representations and are arguably easier to demonstrate due to direct data dependency - ours is the first work to provide concrete evidence and a clear universal hypothesis at the neural parameter level.
Empirical studies further strengthen this perspective. The lottery ticket hypothesis (frankle2019lottery) demonstrates that overparameterized networks contain sparse subnetworks capable of matching full-model performance, implying that task-relevant information resides in a small, structured subset of weights. Similarly, mode connectivity studies (garipov2018mode) reveal that seemingly isolated optima in parameter space are often connected by low-loss paths, suggesting that task solutions lie on a shared manifold. In convolutional models, Krizhevsky et al. (Krizhevsky2012ImageNet) famously observed that early layers consistently learn Gabor-like filters, indicating a universal inductive bias in early representations. More recent works (guth24; guth2024universalityneuralencodingscnns) extends this observation to deeper layers, showing that certain eigenvectors of trained convolutional layers recur across networks trained on different datasets.
While these studies are suggestive of shared structures in neural representations or parameters, they remain limited in their focus, application and analysis. Our work fills this critical gap by presenting a principled and empirically validated method for discovering and utilizing universal parametric subspaces that span across architectures, tasks, and modalities. By conducting large-scale spectral analyses of over large number of diverse architectures, models and tasks, we demonstrate that a small number of principal directions consistently capture the majority of task-relevant variation. We then operationalize these findings by developing a practical framework for reusing these subspaces for parameter-efficient finetuning, task adaptation, and model merging, achieving competitive performance while dramatically reducing memory and compute requirements.
A.2 Theoretical Analysis
We apply a standard generalization bound over the squared error between the task function and its projection onto the shared subspace:
$\ell(f_{t},x)=\|f_{t}(x)-f_{t,k}(x)\|^{2}$
To justify the application of PAC-style bounds, we verify that this loss is bounded. We assume that each task predictor $f_{t}$ lies in a Reproducing Kernel Hilbert Space (RKHS) with norm bounded by $B$, i.e., $\|f_{t}\|_{\mathcal{H}}\leq B$, and that the projection $f_{t,k}$ onto the learned shared subspace $\hat{\mathcal{H}}_{k}$ also satisfies $\|f_{t,k}\|_{\mathcal{H}}\leq B$.
Using the reproducing property and assuming a kernel bound $\kappa^{2}=\sup_{x\in\mathcal{X}}\|\phi(x)\|^{2}$, we have for any $x$:
$\|f_{t}(x)\|\leq\kappa B\quad\text{and}\quad\|f_{t,k}(x)\|\leq\kappa B$
Thus, the pointwise squared loss is bounded as:
$\|f_{t}(x)-f_{t,k}(x)\|^{2}\leq(\|f_{t}(x)\|+\|f_{t,k}(x)\|)^{2}\leq(2\kappa B)^{2}=4\kappa^{2}B^{2}$
Therefore, the loss function is bounded in $0,4\kappa^{2}B^{2}]$, satisfying the conditions required for PAC-style generalization bounds to hold.
Lemma A.1 (Matrix Bernstein for self-adjoint operators).
There exist absolute constants $C>0$ such that, for any $\delta_{T}\in(0,1)$, we have with probability at least $1-\delta_{T}$,
$\big\|\hat{\mathcal{S}}-\mathcal{S}\big\|_{\mathrm{op}}\leq C\,B^{2}\!\left[\sqrt{\frac{\ln(c/\delta_{T})}{T}}\;+\;\frac{\ln(c/\delta_{T})}{T}\right]$
Proof.
Operator Bernstein (intrinsic form).
Let $X_{1},\ldots,X_{T}$ be independent, mean-zero, self-adjoint, bounded operators on a separable Hilbert space. Suppose
$\|X_{t}\|_{\mathrm{op}}\leq L\quad\text{a.s. for all }t.$
Then from (minsker2017extensionsbernsteinsinequalityselfadjoint; koltchinskii2014concentrationinequalitiesmomentbounds) there exist absolute constants $C,c>0$ such that for every $\delta\in(0,1)$,
$\Bigg\|\frac{1}{T}\sum_{t=1}^{T}X_{t}\Bigg\|_{\mathrm{op}}\;\leq\;C\!\left[\sqrt{\,\frac{\Big\|\sum_{t=1}^{T}\mathbb{E}[X_{t}^{2}]\Big\|_{\mathrm{op}}}{T^{2}}\,\ln\!\Bigg(\frac{c\Big(1+\dfrac{\operatorname{tr}\!\big(\sum_{t=1}^{T}\mathbb{E}[X_{t}^{2}]\big)}{\Big\|\sum_{t=1}^{T}\mathbb{E}[X_{t}^{2}]\Big\|_{\mathrm{op}}}\Big)}{\delta_{T}}\Bigg)}\;+\;\frac{L}{T}\,\ln\!\Bigg(\frac{c\Big(1+\dfrac{\operatorname{tr}\!\big(\sum_{t=1}^{T}\mathbb{E}[X_{t}^{2}]\big)}{\Big\|\sum_{t=1}^{T}\mathbb{E}[X_{t}^{2}]\Big\|_{\mathrm{op}}}\Big)}{\delta_{T}}\Bigg)\right]$
with probability at least $1-\delta_{T}$.
Application to $X_{t}=f_{t}^{\star}\otimes f_{t}^{\star}-\mathcal{S}$ with $\|f_{t}^{\star}\|\leq B$ a.s.
We have
$\|X_{t}\|_{\mathrm{op}}\leq\|f_{t}^{\star}\|^{2}+\|\mathcal{S}\|_{\mathrm{op}}\leq B^{2}+\mathbb{E}\|f^{\star}\|^{2}\leq 2B^{2}.$
so $L\leq 2B^{2}$. Moreover, for $X_{t}=f_{t}^{\star}\otimes f_{t}^{\star}-\mathcal{S}$ we have
$\mathbb{E}[X_{t}^{2}]\;\preceq\;2B^{2}\mathcal{S}.$
Hence
[…]
By asumption 2.3,
$\frac{\operatorname{tr}(\sum_{t=1}^{T}\mathbb{E}[X_{t}^{2}])}{\big\|\sum_{t=1}^{T}\mathbb{E}[X_{t}^{2}]\big\|_{\mathrm{op}}}\;\leq\;\frac{\operatorname{tr}(\mathcal{S})}{\|\mathcal{S}\|_{\mathrm{op}}}\;\leq\;\kappa.$
Therefore the intrinsic logarithmic factor in Bernstein reduces to
$\ln\!\left(\frac{c(1+\kappa)}{\delta_{T}}\right),$
and since $\kappa$ is a fixed constant, $1+\kappa$ can be absorbed into $c$.
Plugging into Bernstein gives
$\|\hat{\mathcal{S}}-\mathcal{S}\|_{\mathrm{op}}\;\leq\;C\left[\sqrt{\frac{2B^{2}\|\mathcal{S}\|_{\mathrm{op}}\,\ln(c/\delta_{T})}{T}}\;+\;\frac{2B^{2}\,\ln(c/\delta_{T})}{T}\right],$
with probability at least $1-\delta_{T}$.
Lemma A.2 (Davis–Kahan, sin-$\Theta$).
Let $\gamma_{k}>0$. Then
$\left\lVert\tilde{P}_{k}-P_{k}\right\rVert_{\mathrm{op}}\ \leq\ \frac{2}{\gamma_{k}}\,\left\lVert\tilde{\mathcal{S}}-\mathcal{S}\right\rVert_{\mathrm{op}}.$
using definition of $\gamma_{k}$ from definition 2.1.
Theorem A.3 (Restating Two-level convergence to the shared subspace theorem).
Assume 2.32.4. Let $c_{1},c_{2}$ be any absolute constants. For any $\delta\in(0,1)$, choose $\delta_{t}=\delta/(2T)$ and set $\delta_{T}=\delta/2$. With probability at least $1-\delta$ (over tasks and all per-task samples),
$\left\lVert\tilde{\mathcal{S}}-\mathcal{S}\right\rVert_{\mathrm{op}}\ \leq\ c_{1}B^{2}\sqrt{\frac{\ln(c_{2}/\delta)}{T}}\ +\ (2B\bar{\eta}+\overline{\eta^{2}})$ (3)
If moreover $\gamma_{k}>0$, then
$\left\lVert\tilde{P}_{k}-P_{k}\right\rVert_{\mathrm{op}}\ \leq\ \frac{2}{\gamma_{k}}\!\left(c_{1}B^{2}\sqrt{\frac{\ln(c_{2}/\delta)}{T}}+(2B\bar{\eta}+\overline{\eta^{2}})\right).$ (4)
where $\bar{\eta}=\frac{1}{T}\sum^{T}_{t=1}\eta_{t}$, $\overline{\eta^{2}_{t}}=\frac{1}{T}\sum^{T}_{t=1}\eta^{2}_{t}$ and $\eta_{t}$ is defined same as in assumption 2.4
Proof of Theorem 2.5.
(i) Triangle split. $\left\lVert\tilde{\mathcal{S}}-\mathcal{S}\right\rVert_{\mathrm{op}}\leq\left\lVert\tilde{\mathcal{S}}-\hat{\mathcal{S}}\right\rVert_{\mathrm{op}}+\left\lVert\hat{\mathcal{S}}-\mathcal{S}\right\rVert_{\mathrm{op}}$.
(ii) Within-task term. We know that,
$\left\lVert\hat{f}_{t}\otimes\hat{f}_{t}-f^{\star}_{t}\otimes f^{\star}_{t}\right\rVert_{\mathrm{op}}$ $\leq\left\lVert\hat{f}_{t}-f^{\star}_{t}\right\rVert\,(\left\lVert\hat{f}_{t}\right\rVert+\left\lVert f^{\star}_{t}\right\rVert)$
$\leq\|\hat{f}_{t}-f^{\star}_{t}\|(\|\hat{f}_{t}\|+\|f^{\star}_{t}\|)$
$\leq\eta_{t}\,(2B+\eta_{t})\qquad\text{(since $\|\hat{f}_{t}\|\leq\|f_{t}^{\star}\|+\|\hat{f}_{t}-f_{t}^{\star}\|\leq B+\eta_{t}$)}$
$=2B\,\eta_{t}+\eta_{t}^{2}.$
Averaging and using the triangle inequality for operator norms,
$\left\lVert\tilde{\mathcal{S}}-\hat{\mathcal{S}}\right\rVert_{\mathrm{op}}$ $\leq 2B\,\bar{\eta}+\overline{\eta^{2}}$
This holds on the event $\bigcap_{t=1}^{T}\{\left\lVert\hat{f}_{t}-f_{t}^{\star}\right\rVert\leq\eta_{t}\}$, whose probability is at least $1-\sum_{t}\delta_{t}=1-\delta/2$.
(iii) Across-task term. Let $X_{t}:=f_{t}^{\star}\otimes f_{t}^{\star}-\mathbb{E}[f^{\star}\otimes f^{\star}]$. Then $X_{t}$ are independent, mean-zero, self-adjoint, and $\left\lVert X_{t}\right\rVert_{\mathrm{op}}\leq\left\lVert f_{t}^{\star}\right\rVert^{2}+\left\lVert S\right\rVert_{\mathrm{op}}\leq 2B^{2}$. Lemma A.1 (with $R\asymp B^{2}$) yields
$\left\lVert\hat{\mathcal{S}}-\mathcal{S}\right\rVert_{\mathrm{op}}\leq c_{1}B^{2}\sqrt{\tfrac{\ln(c_{2}/\delta)}{T}}$
$\left\lVert\tilde{\mathcal{S}}-\mathcal{S}\right\rVert_{\mathrm{op}}\$$\leq\ c_{1}B^{2}\sqrt{\frac{\ln(c_{2}/\delta)}{T}}\ +\ 2B\,\left(\sum^{T}_{t=1}\mathcal{R}_{n_{t},D_{t}}(\mathcal{H})\ +\sqrt{\frac{\ln{(2T/\delta)}}{2n_{t}}}\right)+\ \left(\sum^{T}_{t=1}\mathcal{R}^{2}_{n_{t},D_{t}}(\mathcal{H})\ +\frac{\ln{(2T/\delta)}}{2n_{t}}\right)$
$\leq\ c_{1}B^{2}\sqrt{\frac{\ln(c_{2}/\delta)}{T}}\ +\ O\left(\sum^{T}_{t=1}\mathcal{R}^{2}_{n_{t},D_{t}}(\mathcal{H})\ +\frac{\ln{(2T/\delta)}}{2n_{t}}\right)$
with probability at least $1-\delta_{T}=1-\delta/2$.
(iv) Union bound and Davis–Kahan. Combining (ii)–(iii) with a union bound gives equation 1. Lemma A.2 then implies equation 2. ∎
Definition A.4 (Population projection risk).
For a $k$-dimensional subspace $\mathcal{\mathcal{H}}^{\star}_{k}\subset\mathcal{H}$, define
$\mathcal{R}(\mathcal{\mathcal{H}}^{\star}_{k}):=\mathbb{E}_{t\sim\tau}\left\lVert f_{t}^{\star}-P_{\mathcal{\mathcal{H}}^{\star}_{k}}f_{t}^{\star}\right\rVert^{2}.$
Corollary A.5 (Excess projection risk of the learned subspace).
Under the event of Theorem 2.5,
$\mathcal{R}(\tilde{\mathcal{H}}_{k})\ \leq\ \sum_{i>k}\lambda_{i}\ +\ \frac{2\,\operatorname{tr}(S)}{\gamma_{k}}\!\left(c_{1}B^{2}\sqrt{\frac{\ln(c_{2}/\delta)}{T}}+2B\,\bar{\eta}+\overline{\eta^{2}}\right).$
Proof.
Optimality of $P_{k}$ gives $\mathcal{R}(\mathcal{H}_{k}^{\star})=\sum_{i>k}\mu_{i}$. Moreover,
$\mathcal{R}(\tilde{\mathcal{H}}_{k})-\mathcal{R}(\mathcal{H}_{k}^{\star})=\mathbb{E}\left\langle f_{t}^{\star},(P_{k}-\tilde{P}_{k})f_{t}^{\star}\right\rangle\leq\left\lVert\tilde{P}_{k}-P_{k}\right\rVert_{\mathrm{op}}\,\mathbb{E}\left\lVert f_{t}^{\star}\right\rVert^{2}=\operatorname{tr}(S)\,\left\lVert\tilde{P}_{k}-P_{k}\right\rVert_{\mathrm{op}}.$
Apply equation 2. ∎
Remark A.6 (Where Rademacher complexity enters).
Assumption 2.4 is instantiated by your learning procedure. For strongly-convex ERM (e.g., kernel ridge), a standard Rademacher-based excess-risk bound together with curvature yields an $\eta_{t}=\eta_{t}(n_{t},\delta_{t})$ that vanishes with $n_{t}$. Plugging these $\eta_{t}$ into $\bar{\eta}$ and $\overline{\eta^{2}}$ makes the rate explicit.
Appendix B Universal Subspace Analysis
Similar methodology is followed for subspace analysis for both LoRA and classical weight models. In fact, LoRA analysis’ results can be theoretically extended to classical weights, as LoRA weights can be construed to be simple translations from a mean weight matrix. However, in order to solidify our universal subspace hypothesis, we conduct extensive experiments for both types of models. LoRA is chosen because of the recent spurt in the availability of LoRA models trained on diverse kinds of datasets and models. We do this universal subspace analysis on all weight parameters in every neural network layer except the first (or few initial) and last neural network layer. This is because these layers may differ across models due to differences in input shapes and types, loss functions, and the tasks being trained. We also focus our analysis on linear/fully-connected and matrix weights, as the analysis done on these are straightforward and the results observed can be trivially extended to other types of neural parameters (ma2017equivalencefullyconnectedlayer).
Secondary Subspace
refers to the residual subspace that remains after removing the top $k$ principal directions associated with the low-rank universal subspace. This subspace is orthogonal to the universal subspace and serves as a control for evaluating the uniqueness and effectiveness of the learned shared subspace. To make computation tractable when the residual subspace is high-dimensional, we focus on the top components beyond rank $k$, as computing a full SVD is often impractical. This approximation is justified, since the lower components typically capture noise, which has been shown to degrade performance (sharma_laser_2023).
How to choose top $k$ components?
As shown in all eigenvalue (scree) plots, a trivial way to choose is a simple visual inspection, since we can see a discontinuity in the spectral analysis. Another way is to define a threshold on the explained variance, all components whose explained variance is close to zero <.01 are considered secondary subspace, and can be discarded. A more structured way is to define an optimal singular value threshold for the HOSVD, as found by previous works (gavish2014optimalhardthresholdsingular).
B.1 Lower rank shared universal subspaces within low rank adaptation (LoRA) models
Spectral Decomposition is employed to extract the top k principal directions for each of the LoRA matrices B and A, which are concatenated across all available models. Subsequently, the top k principal directions are selected to define the low-rank subspace shared among the LoRA matrices. This process is conducted separately for each layer of the model to derive a low-rank approximated shared subspace for every individual layer. In practice, for every layer, the rank vectors of all available LoRA matrices are extracted and concatenated into a single matrix. This matrix is then normalized by subtracting the feature-wise mean from each vector, after which principal directions are extracted. The mode-1(order-1) variant of our method is mathematically equivalent to Principal Component Analysis (PCA), hence we can use torch.pca_lowrank or sklearn.decomposition.PCA to extract the principal directions. The data matrix corresponding to a specific layer for 500 LoRA models is structured as $500r\times d$, where $r$ denotes the rank of each LoRA and $d$ specifies the dimension of each rank vector. The same calculation can be applied to the BA matrix instead of individually to B and A, thereby increasing the computational cost of the Spectral Decomposition without affecting the outcome.
Figure 8: Spectral analysis of the Mistral-7B-Instruct-v0.2 model: Aggregated eigenvalue (scree) plot across 500 LoRA models and all layers. The plot demonstrates that the majority of the variance is consistently captured by the top 16 principal directions, indicating the presence of a shared low-dimensional universal subspace.
Figure 9: Layerwise Eigenvalue Plots of 500 Mistral-7B-Instruct-v0.2 models. Each layer has 3 sets of parameters - k_proj,q_proj,v_projk\_proj,q\_proj,v\_proj
Universal Mistral-7B/Lots of LoRAs experiment details
In our first experimental analysis, we use 500 LoRA models trained on distinct Natural Instructions (wang-etal-2022-super) using Mistral-7B-Instruct-v0.2 (jiang2023mistral7b) as the base (brüelgabrielsson2024compressserveservingthousands). Please refer to  brüelgabrielsson2024compressserveservingthousands for more details on how the LoRA models were trained.
Table 7: Models from HuggingFace for the Universal Mistral LoRA. Models in blue indicate the OOD models and the ones in red are the IID models used for evaluation.
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task391 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task290
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task442 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1598
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task039
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task076 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task627
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task664 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task819
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1631
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task190 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1391
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1342 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task620
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task769 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1448
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task247 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task513
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task875 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task515
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1534 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1551
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task583 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1431
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task270 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1487
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task679 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task456
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task385 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1607
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task278 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task022
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task210 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task137
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task574 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task629
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1378 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1194
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1529 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task453
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task102 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task460
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1204 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1384
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1572 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task699
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1722 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task580
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task605 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1152
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1283 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task637
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task723 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task084
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task201 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task956
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task167 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1192
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task300 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1714
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task388 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task516
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task127 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task362
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1158 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task322
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task697 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1566
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1451 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1135
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task341 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task267
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1720 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1452
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task131 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task685
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task727 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1590
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1731 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task047
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task929 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1592
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1326 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task615
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1216 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task689
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1156 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1657
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task833 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1206
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1151 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task244
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1562 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task043
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task044 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task722
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task183 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task563
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task155 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task353
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task616 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1724
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task288 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task092
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task707 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task577
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task742 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task706
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1401 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1393
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1198 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task966
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task219 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1211
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task050 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task494
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1379 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task176
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task068 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task566
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task333 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task593
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task667 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1670
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task733 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task472
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1168 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task075
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task148 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task683
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1315 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task121
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task370 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task856
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task891 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task140
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task609 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task344
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1703 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task070
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task072 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task504
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task695 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1434
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task095 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task346
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task274 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1325
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1190 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task568
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1482 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task924
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task761 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task596
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task382 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task926
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task065 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1421
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task323 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1310
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task110 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1288
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1503 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task269
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task821 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task565
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task867 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task755
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task378 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task518
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task195 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task618
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task638 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1217
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task118 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1564
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1429 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task687
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task640 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1328
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1311 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task285
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1341 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1603
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task162 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task063
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task686 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1483
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task146 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1729
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task852 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1622
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1704 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task145
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task648 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task151
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1212 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task045
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task893 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task936
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1425 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1533
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task093 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task389
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task670 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task461
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task113 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task066
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task497 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task343
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task908 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1606
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task381 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task355
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task850 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1557
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task356 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1394
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1428 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task087
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task582 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task770
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task035 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1317
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1203 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1444
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1585 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1508
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task740 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task429
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1427 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task328
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task955 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task130
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task161 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task507
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1502 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task505
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task633 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1645
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1486 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1146
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1380 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1088
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task033 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task085
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1294 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task080
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task489 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1721
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1713 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task721
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1403 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task746
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task728 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task889
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1583 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1665
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task708 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1419
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task963 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task454
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task308 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task828
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task579 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task753
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1404 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1201
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task901 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1567
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1319 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task858
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1200 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task492
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task675 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task094
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1506 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task694
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task614 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1390
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1355 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task714
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task457 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1565
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task834 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task642
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task732 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1605
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task326 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1292
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task716 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1479
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1147 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task153
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1495 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1196
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1489 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task294
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task157 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task147
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1197 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task754
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1599 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1420
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1285 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task587
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1338 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task116
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task713 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task431
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task067 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task934
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task617 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task696
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task846 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task933
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task865 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task671
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1398 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1518
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task628 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1286
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1596 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task298
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task645 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task903
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task594 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task413
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task719 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task672
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1418 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task475
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task357 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1387
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1711 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task304
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task750 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1320
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task034 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task692
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1406 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task119
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task211 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task363
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1087 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task083
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1385 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1308
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1656 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task892
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task499 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1409
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1207 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task564
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task325 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task206
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task890 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1520
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task703 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task318
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task366 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task335
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task600 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task477
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task138 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task291
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task074 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1389
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task192 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task316
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1609 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task296
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task666 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1582
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1728 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task701
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task275 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task107
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task079 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1157
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1167 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task403
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task359 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task517
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task351 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task964
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task904 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1148
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task879 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task636
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1509 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task207
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task228 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task209
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task128 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task710
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1322 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task163
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task178 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task089
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task700 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1581
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task927 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task101
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task123 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1321
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task550 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task129
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task393 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1214
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task277 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1447
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task324 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task455
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task725 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task365
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1316 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1199
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task717 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task245
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task874 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task925
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task380 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1712
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1504 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task619
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task590 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1186
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task736 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task069
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task377 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task181
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task859 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task144
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task632 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task641
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task064 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task630
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1154 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task390
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1188 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task625
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task607 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task495
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1189 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task398
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task108 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1347
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1541 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task202
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1723 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1669
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1089 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1584
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task081 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task329
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task691 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task588
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1593 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task724
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task149 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1449
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1313 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1453
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task905 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task704
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task585 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1209
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task249 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1386
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1400 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task751
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1332 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task674
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task379 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task243
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1318 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task428
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task488 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task705
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task698 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1601
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task861 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1510
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task077 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task509
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task734 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task720
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1210 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task284
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task584 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task105
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task330 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task923
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task319 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task400
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task246 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task726
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1568 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1442
Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task1640 Lots-of-LoRAs/Mistral-7B-Instruct-v0.2-4b-r16-task280
Figure 9 presents the aggregated results across all layers, with error bars representing the standard deviation. For reference, the eigenvalue (scree) plot from  3(b) is also reproduced in  Figure 9. This plot depicts the proportion of variance explained by each principal component, computed across all weight matrices and layers from 500 independently trained Mistral models. The concentration of variance within the top $k$ components reveals the presence of a consistent low-dimensional subspace, offering strong empirical support for the universal subspace hypothesis.
The individual plots provide spectral analysis results for the key, query, and value matrices from all 32 layers of all 500 Mistral models. For clarity, only the top 128 principal directions are visualized, representing a subset of the full component basis. This truncation mitigates the visual distortion caused by the long tail of near-zero eigenvalues beyond the universal subspace, which would otherwise dominate the graph without contributing meaningful information.
[…]
Table 8: Models from HuggingFace used for the Universal Stable Diffusion-XL subspace extraction
alphonse-mucha-style directors-coen-brothers-style larry-carlson-style rene-magritte-style
beeple-mike-winkelmann-style director-sergei-eisenstein-style lascaux richard-corben-style
character-design director-sofia-coppola-style laurel-burch-style richard-dadd-style
director-christopher-nolan-style director-terrence-malick-style lawrence-alma-tadema-style richard-hescox-style
director-lars-von-trier-style director-tim-burton-style leonid-afremov-style richard-scarry-style
director-ridley-scott-style director-wes-anderson-style leonora-carrington-style robert-adams-style
director-stanley-kubrick-style director-wong-kar-wai-style levitating-cube robert-crumb-style
director-zhang-yimou-style director-yorgos-lanthimos-style liam-wong-style robert-rauschenberg-style
olafur-eliasson-style dixit-card-generator lotte-reiniger-style rodney-matthews-style
origami dressed-animals louis-comfort-tiffany-style roger-ballen-style
simone-martini-style dripping-art lovis-corinth-style roger-deakins-style
studio-ghibli-style edward-gorey-style lucas-cranach-style romare-bearden-style
ukiyo-e-art elizabeth-gadd-style luc-schuiten-style ryoji-ikeda-style
wu-guanzhong-style erik-johansson-style lyonel-feininger-style sacha-goldberger-style
1987-action-figure-playset-packaging erik-madigan-heck-style made-of-iridescent-foil salomon-van-ruysdael-style
aardman-animations-style euan-uglow-style makoto-shinkai-style sam-spratt-style
akos-major-style felipe-pantone-style marc-silvestri-style sandy-skoglund-style
albumen-print filip-hodas-style marianna-rothen-style santiago-caruso-style
alec-soth-style folk-art maria-sibylla-merian-style shaun-tan-style
alejandro-jodorowsky-style gabriel-pacheco-style mark-catesby-style shepard-fairey-style
alessandro-gottardo-style gemma-correll-style mark-ryden-style sidney-nolan-style
alex-andreev-style george-condo-style martin-whatson-style simon-stalenhag-style
alex-gross-style gilbert-garcin-style mary-cassatt-style skottie-young-style
alfred-augustus-glendening-style gregory-crewdson-style maurice-de-vlaminck-style sofonisba-anguissola-style
alex-pardee-style gustave-dore-style maurice-prendergast-style sophie-gengembre-anderson-style
alternate-realities hasui-kawase-style maxfield-parrish-style stained-glass-portrait
ando-fuchs-style hiroshi-nagai-style maxime-maufra-style stanley-donwood-style
andre-derain-style infrared-photos mike-mignola-style stephan-martiniere-style
andrei-tarkovsky-style isometric-cutaway mikhail-vrubel-style stephen-gammell-style
andrew-wyeth-style ivan-bilibin-style moebius-jean-giraud-style stop-motion-animation
angus-mckie-style james-c-christensen-style movie-poster surreal-collage
anna-maria-garthwaite-style james-jean-style moving-meditations surreal-harmony
atey-ghailan-style james-r-eads-style nadav-kander-style surreal-plate
audrey-kawasaki-style james-turrell-style natalia-goncharova-style syd-mead-style
avant-garde-fashion jan-brueghel-style n-c-wyeth-style synthwave-t-shirt
banksy-style jan-svankmajer-style needlepoint teamlab-style
bas-relief jan-van-eyck-style neon-night terry-gilliam-style
century-botanical-illustration jan-van-goyen-style nicolas-poussin-style thomas-cole-style
christopher-balaskas-style j-c-leyendecker-style noah-bradley-style thomas-kinkade-style
christopher-ryan-mckenney-style jean-baptiste-camille-corot-style ohara-koson-style thomas-moran-style
clay-animation jean-baptiste-monge-style okuda-san-miguel-style thomas-schaller-style
color-palette jean-baptiste-simeon-chardin-style olly-moss-style tim-walker-style
craig-mullins-style jean-metzinger-style op-art tintoretto-style
crocheted jean-michel-basquiat-style parralel-dimensions todd-hido-style
daniel-arsham-style jessie-willcox-smith-style pascal-campion-style tove-jansson-style
dark-fantasy jim-mahfood-style paul-gustav-fischer-style tracie-grimwood-style
dave-mckean-style john-albert-bauer-style paul-laffoley-style vasily-vereshchagin-style
diorama john-berkey-style paul-signac-style vertical-landscapes
director-agnes-varda-style john-blanche-style peter-doig-style victor-brauner-style
death-stranding john-constable-style peter-paul-rubens-style victor-moscoso-style
director-akira-kurosawa-style john-everett-millais-style philippe-druillet-style video-installation
director-andrei-zvyagintsev-style john-harris-style photographer-elena-helfrecht-style vintage-postage-stamps
director-bong-joon-ho-style john-james-audubon-style photographer-flora-borsi-style weegee-style
director-darren-aronofsky-style john-kenn-mortensen-style photographer-maren-klemp-style wendy-froud-style
director-david-fincher-style john-martin-style photographer-martin-kimbell-style will-eisner-style
director-david-lynch-style john-singer-sargent-style photographer-reuben-wu-style willem-haenraets-style
cute-animals john-singleton-copley-style pierre-auguste-renoir-style willem-van-aelst-style
ben-aronson-style john-william-waterhouse-style pierre-bonnard-style william-langson-lathrop-style
director-emir-kusturica-style joseph-wright-of-derby-style pieter-claesz-style william-mctaggart-style
director-gaspar-noe-style josh-agle-style punk-collage william-merritt-chase-style
director-jean-pierre-jeunet-style josh-kirby-style quentin-blake-style winslow-homer-style
director-krzysztof-kieslowski-style jules-bastien-lepage-style raimonds-staprans-style worthington-whittredge-style
director-martin-scorsese-style kate-greenaway-style ralph-bakshi-style yaacov-agam-style
director-nicolas-winding-refn-style kay-nielsen-style ralph-steadman-style yoh-nagao-style
director-park-chan-wook-style kilian-eng-style randolph-caldecott-style yves-klein-style
director-pedro-almodovar-style kirigami ray-caesar-style zanele-muholi-style
director-quentin-tarantino-style konstantin-korovin-style remedios-varo-style
Universal SDXL experiment details
Our second experiment involves the complex and multimodal task of Text-to-Image generation using the Stable Diffusion-XL model sdxl. We extract our low rank universal subspace from publicly available LoRA models on HuggingFace repository von-platen-etal-2022-diffusers - LABEL:tab:diffusion_data lists all the SDXL models that we used to extract the Universal Subspace. As can be seen in LABEL:tab:diffusion_data, the models range wildly in styles on which they were finetuned. The fact that all these diverse models can be represented by a single low rank universal subspace model strongly verifies our hypothesis. We use top 16 components and 30 denoising steps. For each experiment model shown in Table 1 and Figure 5, that LoRA model is reconstructed using a universal subspace created using rest of the available LoRA adapters, essentially confirming the generalization capability of this subspace.
We then use this single SDXL universal subspace to generate images with similar styles to evaluate whether this subspace is capable of doing so, by projecting randomly chosen LoRA models into this subspace. Figure 5 shows that our universal subspace matches the visual quality and style nuances of individual LoRAs, resulting in significant memory savings. Table 1 shows quantitative results for our Universal subspace in terms of CLIP scores, where interestingly we can see that our Universal Subspace outperforms the individual LoRA models. This improvement may be attributed to our Universal SDXL removing noise from the subspace - a phenomenon previously observed by sharma_laser_2023. The styles used in Table˜1, which are in LABEL:tab:diffusion_data are (from Style 1 to Style 10) Ukiyo-e Style, Todd Hildo Style , Olly Moss Style , Needlepoint Style , Studio Ghibli Style, Surreal Harmony Style , Dressed Animal Style , Lascaux Cave Art Style , Kirigami Style , Yaacov Agam Style.
B.2 Low rank shared universal subspaces in classical weights
Figure 10: Spectral analysis of the Vision Transformer (ViT-base-patch16-224) model: Aggregated eigenvalue (scree) plot across 500~500 ViT models and all layers. The plot demonstrates that the majority of the variance is consistently captured by the top 16 principal directions, indicating the presence of a shared low-dimensional universal subspace.
In order to further solidify the evidence for our universal subspace hypothesis, we show that this universality does extend beyond adapter models to conventional weights. We do not focus on convolutional weight parameters as they can simply be equated with fully connected layers (ma2017equivalencefullyconnectedlayer), and have been shown, in limited scope, to match Gabor-like filters (Krizhevsky2012ImageNet). Therefore, our analysis trivially extends to these kinds of parameters as well. However, there are a few practical differences between the low rank adapter and classical weight subspace analysis. The classical weight subspace analysis is more computationally expensive relative to the LoRA one due to high dimensionality of the parameters, but in effect, same. Additionally, the number of sufficiently well trained models is understandably fewer than LoRA models. Further, there is also higher variance in terms of model quality in the classical weights as it is harder to optimize these models as compared to LoRA which often are optimized from a good initialization point (the pretrained base model). An outcome of this is that the universal subspace approximation that we obtain from the publicly available pretrained models are noisier than their LoRA counterparts. Inspite of this, our universal subspace hypothesis remains validated.
To further support our universal subspace hypothesis, we extend our analysis beyond adapter models to standard full-rank weights. We exclude convolutional parameters from explicit consideration, as they are functionally equivalent to fully connected layers under certain conditions (ma2017equivalencefullyconnectedlayer), and their learned representations (e.g., Gabor-like filters) have been studied, in limited scope, in prior work (Krizhevsky2012ImageNet). Consequently, our analysis generalizes naturally to convolutional weights as well.
There are, however, practical differences between the subspace analysis of full-rank model weights and that of low-rank adapters. First, analyzing conventional weight matrices is significantly more computationally intensive because of their higher dimensionality. Second, the availability of a large number of independently and sufficiently well-trained models is more limited compared to LoRA models. Third, the classical weight models exhibit greater variance in model quality, since they must be trained from scratch, often without the benefit of a well-optimized initialization, unlike LoRA which builds upon a strong pretrained base.
As a result, the subspaces estimated from classical weights tend to be noisier, and the universality signal is less pronounced. Despite these challenges, we still observe consistent structure in the leading components, lending further empirical support to the universal subspace hypothesis.
Figure 11: Layerwise Eigenvalue Plots of 500 ViT models.
Universal ViT-base-patch16-224 experiment details
We collect $\sim$500 pretrained ViT models from HuggingFace, shown in LABEL:tab:vit_models, spanning very diverse domains — many of which would be considered orthogonal to one another in terms of domain generalization. These models have been trained with varying losses, optimizers, and initializations. These models were used as-is, without curation or access to training data, to reflect real-world variability. Figure 10 shows the summarized scree plot for all relevant layers of ViT (sans first and last layers due to differences in shape and tasks) for all $\sim 500$ ViT models showing that the majority of variance is captured by the top 16 principal directions, revealing a highly compressible, shared subspace across layers. Only the top 100 components are visualized for clarity, although the available subspace is significantly larger, underlying the sparsity of this universal subspace. We observe this for layerwise analysis in Figure 11 as well. For the experimental results presented in Table 3, we randomly choose 4-5 IID and 4-5 OOD models from LABEL:tab:vit_models for which evaluation dataset is available, and reconstruct these model weights by projecting them into our 16 component universal subspace. For the OOD case, we ensure that the models being evaluated are not present in the subset used for creating the universal subspace approximation. As seen from the results, our extremely sparse subspace model performs competitively compared to the fully trained versions. It is likely that with more careful choice of principal directions per layer would allow for at par or even better performance.
Table 9: Finetuned Models from HuggingFace used for the Universal Vision Transformer subspace extraction (vit-base-patch16-224)
0.50-200Train-100Test-vit-base 2025-01-21-16-13-04-vit-base-patch16-224
2025-02-05-14-22-36-vit-base-patch16-224 21BAI1229
Accomodation_room_classification adam_VitB-p16-224-1e-4-batch_16_epoch_4_classes_24
age_face_detection_base AIvisionGuard-v2
alea amns
AnimeCharacterClassifierMark1 autotrain-48ci8-roib9
autotrain-8oqr6-image0807-20 autotrain-ap-pass-fail-v1
autotrain-g2g80-iwcfm autotrain-google-vit-13epoch
autotrain-ht4es-gbvmt autotrain-image-classifier-cats-and-dogs
autotrain-pknu0-o76h9 autotrain-s0sds-erede
autotrain-test-image-classification autotrain-vit-base-patch16-224-fog-or-smog-classification
beauty-ornot beer-classifier
bg-classif bigger-chord-finetuned
brain-tumor-44 ButterflyClasifModel
camera-type Caracam
cards-vit-base-patch16-224-finetuned-v1 carmodel
cats123 cats-dogs-2024
cats-dogs-classification CheXpert-ViT-U-MultiClass
CheXpert-ViT-U-SelfTrained chord-final-model
chord_ViT-finetuned cifar10-lt
city_multiclass_classification clasificador_masas
corals_binary_classification custom
detect_meme dog-breeds-classification
dog-cat-demo-20240815 dog-cats-model
dummy_classification_model dvm-cars-vit-first-5k
ecg-image-multilabel-classification emotion
EmotionAgeModel emotion_model
emotion-recognition emotion_recognition
emotion_recognition_results emotion-vit
face_age_detection_base_v2 face_age_detection_base_v3_weighted
final-run finetune-cats
fine-tuned finetuned-amazon
fine-tuned-augmented finetuned-bin
finetuned-cifar10 finetuned-indian-food
fine-tuned-model finetuned_model
Fine-Tuned_Model Fine-Tuned_Model2
Fine-Tuned_Model3 Fine-Tuned_Model3_Transfer_learning
finetune-vit-base-patch16-224 finetune_vit_base_patch16_224_1epoch
Flowers food
food-101-finetuned-model Freshness-Fruit_Vegies
frost-vision-v2-google_vit-base-patch16-224 frost-vision-v2-google_vit-base-patch16-224-v2024-11-09
frost-vision-v2-google_vit-base-patch16-224-v2024-11-11 frost-vision-v2-google_vit-base-patch16-224-v2024-11-14
fruit_classification fruits-360-16-7
ft_stable_diffusion gender
giecom-vit-model-clasification-waste google-vit-base-patch16-224-batch32-lr0.0005-standford-dogs
google-vit-base-patch16-224-batch32-lr0.005-standford-dogs google-vit-base-patch16-224-batch32-lr5e-05-standford-dogs
google-vit-base-patch16-224-batch64-lr0.005-standford-dogs google-vit-base-patch16-224-OrganicAndInorganicWaste-classification
google-vit-base-patch16-224-Waste-O-I-classification hf_vit_format_hap_pretrained_256_128
Human-Action-Recognition-VIT-Base-patch16-224 human-actions
image-classification image_classification
image_strawbery-peach_classifier isa-vit_model
lixg_food_model001 Maggi-Parle-G_Classifier
mammals_multiclass_classification MemeDetector
model Model
model-vit-base-finetuned MRI_vit
my_chest_xray_model myclass
my_classification MyPetModel
out outputs
PagesClassificationModel physiotheraphy-E2
plant_disease_detection-beans pokemon_classification
pokemon_model pokemon-vit
recaptcha recycled_waste_classification
results rmsprop_VitB-p16-224-1e-4-batch_16_epoch_4_classes_24
rmsprop_VitB-p16-224-2e-4-batch_16_epoch_4_classes_24 road-conditions
rose_recognition rotated2
Ruster S1_M1_R1_vit_42498800
S1_M1_R1_vit_42509509 S1_M1_R1_ViT_42616100
S1_M1_R2_vit_42498972 S1_M1_R2_ViT_42618476
S1_M1_R3_vit_42499444 S1_M1_R3_ViT_42618486
S2_M1_R1_vit_42499480 S2_M1_R1_ViT_42618522
S2_M1_R2_vit_42499499 S2_M1_R2_ViT_42618530
S2_M1_R3_vit_42499514 S2_M1_R3_ViT_42618549
S5_M1_fold1_vit_42499955 S5_M1_fold1_ViT_42618571
S5_M1_fold2_vit_42499968 S5_M1_fold2_ViT_42618583
S5_M1_fold3_vit_42499983 S5_M1_fold3_ViT_42618589
S5_M1_fold4_vit_42499997 S5_M1_fold4_ViT_42618593
S5_M1_fold5_vit_42500027 S5_M1_fold5_ViT_42621111
Screenshots_detection_to_classification sign-lan-model
square_run_32_batch square_run_age_gender
square_run_first_vote_full_pic_50 square_run_first_vote_full_pic_50_age_gender
square_run_first_vote_full_pic_75 square_run_first_vote_full_pic_75_age_gender
square_run_second_vote square_run_second_vote_full_pic_50
square_run_second_vote_full_pic_50_age_gender square_run_second_vote_full_pic_75
square_run_second_vote_full_pic_75_age_gender square_run_second_vote_full_pic_age_gender
square_run_second_vote_full_pic_stratified square_run_square_run_first_vote_full_pic_25
square_run_square_run_first_vote_full_pic_25_age square_run_square_run_first_vote_full_pic_25_age_gender
square_run_square_run_first_vote_full_pic_25_age_gender_double_check square_run_square_run_second_vote_full_pic_25
square_run_square_run_second_vote_full_pic_25_age_gender square_run_with_16_batch_size
square_run_with_actual_16_batch_size stool-condition-classification
swaddling-classifier swin-tiny-patch4-window7-224-finetuned-eurosat-kornia
tarread telidermai
test-cifar-10 traffic-levels-image-classification
Train-Augmentation-vit-base trainer_output
Train-Test-Augmentation-V3D-vit-base UL_base_classification
UL_bedroom_classification UL_exterior_classification
UL_interior_classification vehicle_multiclass_classification
ViT_ASVspoof_DF vit-augmentation
vit-b16-plant_village vit_base
vit-base-1e-4-15ep vit-base-1e-4-20ep
vit-base-1e-4-randaug vit-base-1stGen-Pokemon-Images
vit-base-25ep Vit-Base-30VN
vit-base-3e-5-randaug vit-base-5e-4
vit-base-add-2-decay vit-base-augment
vit-base-batch-32 vit-base-beans
vit-base-brain-mri vit-base-cat_or_dog
vit-base-change-arg vit-base-cocoa
ViT-Base-Document-Classifier vit-base-fashion
vit-base-finetuned-cephalometric vit-base-food101
vit-base-fruits-360 vit-base-hate-meme
vit-base-nationality vit-base-org-plot
vit-base-oxford-brain-tumor vit-base-oxford-brain-tumor_try_stuff
vit-base-oxford-brain-tumor_x-ray vit-base-oxford-iiit-pets
vit-base-oxford-pets-krasuluk vit-base-patch16-224
vit-base-patch16-224-13_model vit-base-patch16-224-30-vit
vit-base-patch16-224-9models vit-base-patch16-224-abhi1-finetuned
vit-base-patch16-224_augmented-v2_fft vit-base-patch16-224_augmented-v2_tl
vit-base-patch16-224-blur_vs_clean vit-base-patch16-224-brand
vit-base-patch16-224-classifier vit-base-patch16-224-clothes-filter
vit-base-patch16-224-cl-v1 vit-base-patch16-224-crochets-clothes-classification
vit-base-patch16-224-Diastar vit-base-patch16-224-Diastarallclasses
vit-base-patch16-224-dmae-va-U vit-base-patch16-224-dmae-va-U5-100-iN
vit-base-patch16-224-dmae-va-U5-10-45-5e-05 vit-base-patch16-224-dmae-va-U5-20-45-5e-05
vit-base-patch16-224-dmae-va-U5-40-45-5e-05 vit-base-patch16-224-dmae-va-U5-42B
vit-base-patch16-224-dmae-va-U5-42C vit-base-patch16-224-dmae-va-U5-42D
vit-base-patch16-224-ethos vit-base-patch16-224-ethos-25
vit-base-patch16-224-ethos-8 vit-base-patch16-224-ethos-data
vit-base-patch16-224-ethosrealdata vit-base-patch16-224-fatigue
vit-base-patch16-224-finalterm vit-base-patch16-224-finetuned
vit-base-patch16-224-finetuned-barkley vit-base-patch16-224-finetuned-brain-tumor-classification
vit-base-patch16-224-finetuned-Brain-Tumor-Classification vit-base-patch16-224-finetuned-cassava-leaf-disease
vit-base-patch16-224-finetuned-cedar vit-base-patch16-224-finetuned-cifar10
vit-base-patch16-224-finetuned-combinedSpiders vit-base-patch16-224-finetuned-context-classifier
vit-base-patch16-224-finetuned-covid_ct_set_full vit-base-patch16-224-finetuned-covid_ct_set_resumed
vit-base-patch16-224-finetuned-crochets-clothes vit-base-patch16-224-finetuned-dangerousSpiders
vit-base-patch16-224-finetuned-eurosat vit-base-patch16-224-finetuned-feature-maps-v3
vit-base-patch16-224-finetuned-feature-map-v2 vit-base-patch16-224-finetuned-fibre
vit-base-patch16-224-finetuned-flower vit-base-patch16-224-finetuned-flower-classify
vit-base-patch16-224-finetuned-flowers vit-base-patch16-224-finetuned-food101
vit-base-patch16-224-finetuned-food102 vit-base-patch16-224-finetuned-foveated-features
vit-base-patch16-224-finetuned-foveated-features-v2 vit-base-patch16-224-finetuned-galaxy10-decals
vit-base-patch16-224-finetuned-hateful-meme-restructured vit-base-patch16-224-finetuned-hateful-meme-restructured-balanced
vit-base-patch16-224-finetuned-imagegpt vit-base-patch16-224-finetuned-ind-17-imbalanced-aadhaarmask
vit-base-patch16-224-finetuned-ind-17-imbalanced-aadhaarmask-new-parameter vit-base-patch16-224-finetuned-landscape-test
vit-base-patch16-224-finetuned-lora-oxford-pets vit-base-patch16-224-finetuned-masked-hateful-meme-restructured
vit-base-patch16-224-finetuned-noh vit-base-patch16-224-finetuned-original-images
vit-base-patch16-224-finetuned-pneumonia-detection vit-base-patch16-224-finetuned-polyterrasse
vit-base-patch16-224-finetuned-skin vit_base_patch16_224-finetuned-SkinDisease
vit-base-patch16-224-finetuned-teeth_dataset vit-base-patch16-224-finetuned-trash-classifications-albumentations
vit-base-patch16-224-finetuned-turquoise vit-base-patch16-224-finetuned-Visual-Emotional
vit-base-patch16-224-finetuned-vit vit-base-patch16-224-finetune_test
vit-base-patch16-224-food101-16-7 vit-base-patch16-224-food101-24-12
vit-base-patch16-224-for-pre_evaluation vit-base-patch16-224-fruits-360-16-7
vit-base-patch16-224-high-vit vit-base-patch16-224-jvadlamudi2
vit-base-patch16-224-masaratti vit-base-patch16-224-mascotas
vit-base-patch16-224-mascotas-DA vit-base-patch16-224-MSC-dmae
vit-base-patch16-224-newly-trained vit-base-patch16-224-oxford-pets-classification
vit-base-patch16-224-perros-y-gatos vit-base-patch16-224-pure-ViT
vit-base-patch16-224-R1-10 vit-base-patch16-224-R1-40
vit-base-patch16-224-Rado_5 vit-base-patch16-224_rice-disease-02
vit-base-patch16-224_rice-leaf-disease-augmented_fft vit-base-patch16-224_rice-leaf-disease-augmented_tl
vit-base-patch16-224_rice-leaf-disease-augmented-v4_fft vit-base-patch16-224_rice-leaf-disease-augmented-v4_tl
vit-base-patch16-224_rice-leaf-disease-augmented-v4_v5_fft vit-base-patch16-224_rice-leaf-disease-augmented-v4_v5_pft
vit-base-patch16-224-rotated-dungeons-v101 vit-base-patch16-224-rotated-dungeons-v103
vit-base-patch16-224-RU2-10 vit-base-patch16-224-RU2-40
vit-base-patch16-224-RU3-10 vit-base-patch16-224-RU3-40
vit-base-patch16-224-RU4-10 vit-base-patch16-224-RU4-40
vit-base-patch16-224-RU5-10 vit-base-patch16-224-RU5-10-8
vit-base-patch16-224-RU5-40 vit-base-patch16-224-RU9-24
vit-base-patch16-224-RX1-24 vit-base-patch16-224-RX2-12
vit-base-patch16-224-RXL1-24 vit-base-patch16-224-type
vit-base-patch16-224-U6-10 vit-base-patch16-224-U7-10
vit-base-patch16-224-U8-10 vit-base-patch16-224-U8-10b
vit-base-patch16-224-U8-10c vit-base-patch16-224-U8-40
vit-base-patch16-224-U8-40b vit-base-patch16-224-U8-40c
vit-base-patch16-224-U8-40d vit-base-patch16-224-ve-b-U10-12
vit-base-patch16-224-ve-b-U10-24 vit-base-patch16-224-ve-b-U10-40
vit-base-patch16-224-ve-U10-12 vit-base-patch16-224-ve-U10-24
vit-base-patch16-224-ve-U11-12 vit-base-patch16-224-ve-U11-b-24
vit-base-patch16-224-ve-U11-b-40 vit-base-patch16-224-ve-U11-b-80
vit-base-patch16-224-ve-U12-b-24 vit-base-patch16-224-ve-U12-b-80
vit-base-patch16-224-ve-U13-b-120 vit-base-patch16-224-ve-U13-b-24
vit-base-patch16-224-ve-U13-b-80 vit-base-patch16-224-ve-U13b-80R
vit-base-patch16-224-ve-U13b-80RX vit-base-patch16-224-ve-U13b-80RX1
vit-base-patch16-224-ve-U13b-80RX3 vit-base-patch16-224-ve-U13b-R
vit-base-patch16-224-ve-U14-b-24 vit-base-patch16-224-ve-U15-b-80
vit-base-patch16-224-ve-U16-b-80 vit-base-patch16-224-ve-Ub
vit-base-patch16-224-vit vit-base-patch16-224-vit-base-patch16-224-vit-base-patch16-224-dogORnot
vit-base-pets vit-base-PICAI
vit-base-seed-1e-4 vit-base-seed-3e-4
vit-base-travel-document-classification vit-base-v1-eval-epoch-maxgrad-decay-cosine
vit-beans-classifier vit-beta1-0.85
vit-beta1-0.88 vit-beta1-0.95
vit-beta2-0.99 vit-beta2-0.995
vit-beta2-0.9995 vit-bird
ViT_bloodmnist ViT_bloodmnist_std_0
ViT_bloodmnist_std_15 ViT_bloodmnist_std_30
ViT_bloodmnist_std_45 ViT_bloodmnist_std_60
ViT_breastmnist ViT_breastmnist_std_0
ViT_breastmnist_std_15 ViT_breastmnist_std_30
ViT_breastmnist_std_45 ViT_breastmnist_std_60
VIT-cats-vs-dogs vit-cifar10-fine-tuned
vit-class-weight vit-cxr4
vit-demo ViT_dog_food
vit-dropout-0.2 vit-dropout-0.3
vit-dropout-0.4 vit-dropout-0.5
vit-ds-processed vit-emotion-model
vit-epsilon-1e-7 vit-epsilon-1e-9
vit-epsilon-5e-9 vit-face-project-piyush
vit-fine-tune-classification-cats-vs-dogs vit-finetuned-1
vit-food-classification-chrisis2 vit-geometric-shapes-base
vit-google-model-30-classes vit_google_vehicle_classification_model
vit-historical-page vit_Liveness_detection_v1.0
vit-lr-0.0001 vit-lr-0.001
vit-lr-0.01 vit-lr-cosine-restarts
vit-lr-cosine-warm-restarts vit-lr-cosine-warmup
vit-lr-exponential vit-lr-inverse-sqrt
vit-lr-linear vit-lr-poly
vit-lr-reduce-plateau vit-lr-step
vit-mae-base-finetuned-eurosat vit-molecul
vit-ori-dataset-exp vit-plant-classification
vit-plantnet300k vit-plants
vit-real-fake-classification-v1 vit-real-fake-classification-v2
vit-real-fake-classification-v3 vit-real-fake-classification-v4
vit-skin-demo-v1 vit-skin-demo-v2
vit-skin-demo-v3 vit-skin-demo-v4
vit-skin-demo-v5 vit-spam
vit-sports-cls vit-transfer-learning
vit_transformer_eye_disease vit_tumor_classifier
vit-vit vit-vit-base-patch16-224-finetuned-chest-xray
vit-weight-decay-1e-2 vit-weight-decay-1e-3
vit-weight-decay-1e-4 vit-weight-decay-1e-5
wmc_v2_vit_base_wm811k_cls_contra_learning_0916 wmc_v2_vit_base_wm811k_cls_contra_learning_0916_9cls
wmc-wmk811-v0-vit-special_map_det_0917 WS800_ViT_42820348
WS800_ViT_42895082 xraynewww
yet-another-amber-mines zdravJEM_CV_BERT
Universal LLaMA3-8B Experiment Details
To further stress-test our universal subspace hypothesis on classical weight matrices, we extract a shared subspace from approximately 50 finetuned LLaMA3 models, each with 8 billion parameters. These models were obtained from publicly available repositories on HuggingFace. Due to their scale, we do not apply any model selection or filtering, and instead include the entire available set.
As shown in Figure 12, which presents the aggregated scree plot across all layers and all 50 models, the principal variance is concentrated in the top few components—consistent with the emergence of a low-rank universal subspace. For reference, the plot displays only the top 300 components, which represent a small fraction of the full rank, highlighting the inherently low-dimensional structure.
The models included in this analysis span a diverse range of domains, including medical applications, multilingual dialogue systems, and general-purpose assistants, as listed in Table 10. To the best of our knowledge, this is the first work to demonstrate that such a large and heterogeneous collection of high-capacity language models can be jointly represented within a single low-rank subspace.
The layerwise spectral analysis, shown in Figure 13, corroborates this finding: across all layers, the majority of eigenvalues fall below a threshold of <0.001<0.001, indicating that most directions in parameter space contribute negligibly to variation across models. The plots are cropped to show only the leading components due to the large number of total dimensions. We recommend zooming in for clearer visualization.
Figure 12: Spectral analysis of 50 LLaMA-3-8B model: Aggregated eigenvalue (scree) plot across $50$ LlaMa-8B models and all layers. The plot demonstrates that the majority of the variance is consistently captured by few top principal directions, indicating the presence of a shared low-dimensional universal subspace.
Figure 13: Layerwise Scree Plots for 50 LLaMA-3-8B Models. For enhanced clarity, each subplot presents a truncated view of the total possible principal directions. These plots consistently demonstrate that the dominant information, as represented by explained variance, resides within a small number of leading principal directions for all models. Components beyond this initial set are characterized by eigenvalues approaching zero, signifying their redundancy for the universal subspace.
Table 10: Models from HuggingFace used for the Universal LlaMa3-8B subspace extraction
Meta-Llama-3-8B-Instruct-Jailbroken Llama-3-13B-Instruct large_crafting_sft_success suzume-llama-3-8B-multilingual
summary-llama3-8b-f16-full Llama-3-13B-Instruct-v0.1 Llama-3-8B-ProLong-64k-Base LLaMAntino-3-ANITA-8B-Inst-DPO-ITA
ai-medical-model-32bit filtered_crafting_train_data_shorter_length Llama-3-portuguese-Tom-cat-8b-instruct Llama-3-MAAL-8B-Instruct-v0.1
Human-Like-LLama3-8B-Instruct LLaMA-3-8B-Instruct-TR-DPO CabraLlama3-8b chartgpt-llama3
KoLlama-3-8B-Instruct honeypot-llama3-8B Llama-SEA-LION-v2-8B TR
Llama3-8B-Instruct-Turkish-Finetuned Llama-3-15B-Instruct-zeroed Llama-3-8B-Instruct-TAR-Bio-v2 Bio-Medical-Llama-3-8B
filtered_construction_train_data shisa-v1-llama3-8b REFUEL-Llama-3-Armo-iter_1 llama3-instrucTrans-enko-8b
Llama-3-8B-Instruct-Ja llama3-passthrough-chat RoLlama3-8b-Instruct Lloro-SQL
Summary_L3_1000steps_1e7rate_SFT2 CyberSentinel Meta-Llama-3-8B-Instruct-function-calling-json-mode MARS
Llama-3-8B-Instruct-Finance-RAG LLaMA3-Instruct-8B-FR-Spec Llama-3-8B-Japanese-Instruct Llama3-8B-Chinese-Chat
llama-3-chinese-8b-instruct-v2 Athene-RM-8B Llama-3-OffsetBias-RM-8B large_cooking_sft_success
suzume-llama-3-8B-japanese llama-3-chinese-8b-instruct-v3 Waktaverse-Llama-3-KO-8B-Instruct llama-3-8b-gpt-4o-ru1.0
Llama-3-Aplite-Instruct-4x8B-MoE Llama-3-8B-Instruct-DPO-v0.3
Universal Flan-T5 Experiment Details
We collected Flan-T5 models fine-tuned on individual datasets from the GLUE (glue) benchmark. We extract the joint subspace from these models and trends similar to those observed above are seen. This shows that across diverse datasets and tasks a low-rank subspace emerges.
Table 11: Finetuned Flan-T5 Models from HuggingFace used for the Universal Flan-T5 subspace extraction
tanganke/flan-t5-base_glue-cola tanganke/flan-t5-base_glue-mnli
tanganke/flan-t5-base_glue-mrpc tanganke/flan-t5-base_glue-qnli
tanganke/flan-t5-base_glue-rte tanganke/flan-t5-base_glue-qqp
tanganke/flan-t5-base_glue-sst2 tanganke/flan-t5-base_glue-stsb
B.3 Ablating number of models and subspace effectiveness
Although this is implicitly addressed through our large-scale experiments (500 ViTs, 500 Mistral-7B and 300 Stable Diffusion LoRAs, 50 LLaMA3-8B, 177 GPT-2s, Flan-T5, and ResNet50 models) in all Figures and Tables, which demonstrate consistent behavior at different scales. Theorem˜2.5 provides insights on the saturation dynamics where we see that the rate of convergence of the shared subspace to the true subspace is in the order $O(1/T)$, where T is the number of tasks, indicating increasingly effective coverage as T increases. In practice, the minimum number of models per architecture needed to achieve saturation point depends on the quality of the trained models, the diversity of data they have been trained on, and on the architecture itself. Ablating these would require access to all the data for all the models, and very careful training on every training for each data, and then running permutation with all possible combinations of models. All of this is out of reach for most researchers simply due to time, data and compute constraints. We, however, do provide an initial ablation here. For LoRA models shown in LABEL:tab:mistral_models, we choose 9 random (OOD) tasks (39 ,190 ,280 ,290 ,391 ,442 ,1342 ,1391 ,1598) and extract the Universal Subspace from rest of the the tasks, sampled randomly for increasing number of models. The coefficients for OOD tasks are analytically reconstructed to effectively evaluate the universal subspace created from varying number of models. Table˜12 shows that the adequate principal components are quickly extracted, and increasing the number of models has diminishing returns.
Table 12: Lots of LoRAs (Mistral-7B) OOD evaluation per increasing number of models used to extract Universal Subspace
Method Model Number Rouge-L Score
Normal Model - 73.7
Universal model 50 55.8
Universal model 150 66.1
Universal model 250 71.9
Universal model 450 72.3
Appendix C Finding universal subspaces and applying them to future tasks
In this section, we present two tasks, GLUE (glue) and Image Classification. For each experiment, the joint subspace is created using all other models in subset. For Image Classification, we use $k=4$ and train only 8 epochs using learning rate of 1e-4. Importantly, only the coefficients are trained for the experiment. It is important to note that our shared subspace model performs quite well despite using very few (4-5) models to extract the subspace. For GLUE, we use 16-32 components for our subspace, with learning rate of 4e-4, batch size of 64, and 30-80 epochs for each task. In addition, it is likely that our model might perform similarly or better if trained longer or with optimized hyperparameters.
Compute Resources
We conduct all our experiments using a single A5000 GPU, and a CPU with 8 workers. For the universal subspace extraction, all calculation can be done on the CPU. However, GPU would increase the speed of calculation as the layerwise subspace extraction can be parallelized.
Appendix D Discussion and Broader Impact
Our findings suggest that deep neural networks trained across diverse tasks and modalities systematically converge to shared, low-dimensional subspaces within their parameter space. The existence of such universal subspaces challenges conventional assumptions about the independence and diversity of model and task-specific finetuning trajectories. Instead, it highlights a powerful regularity in the way deep models encode task-specific knowledge - one that can be exploited for significantly improved training and deployment efficiency. By leveraging these subspaces, we demonstrate that models can be adapted to new tasks by learning only a small number of coefficients, rather than retraining or storing full sets of weights. This facilitates more robust multi-task learning, model merging, and scalable fine-tuning, with theoretical guarantees and empirical validation across multiple architectures.
The broader societal impact of this work is substantial. Our approach enables large-scale models to be reused and extended with dramatically reduced computational overhead, addressing both the financial and environmental costs associated with training and deploying deep learning systems. This contributes directly to the goals of sustainable and accessible AI. By lowering the hardware and energy requirements for adaptation and inference, we empower under-resourced researchers, institutions, and communities to build upon state-of-the-art models without needing extensive compute infrastructure. Furthermore, by supporting modular model design and data-free model merging, our work lays the foundation for more interpretable, maintainable, and equitable AI systems.